### v2

# Convergence Guarantees of Model-free Policy Gradient Methods for LQR with Stochastic Data

Abstract
Policy gradient (PG) methods are the backbone of many reinforcement learning algorithms due to their good performance in policy optimization problems. As a gradient-based approach, PG methods typically rely on knowledge of the system dynamics. If this is not available, trajectory data can be utilized to approximate first-order information. When the data are noisy, gradient estimates become inaccurate and a study that investigates uncertainty estimation and the analysis of its propagation through the algorithm is currently missing. To address this, our work focuses on the Linear Quadratic Regulator (LQR) problem for systems subject to additive stochastic noise. After briefly summarizing the state of the art for cases with a known model, we focus on scenarios where the system dynamics are unknown, and approximate gradient information is obtained using zeroth-order optimization techniques. We analyze the theoretical properties by computing the error in the estimated gradient and examining how this error affects the convergence of PG algorithms. Additionally, we provide global convergence guarantees for various versions of PG methods, including those employing adaptive step sizes and variance reduction techniques, which help increase the convergence rate and reduce sample complexity. This study contributed to characterizing robustness of the study of the robustness of model-free PG methods, aiming to identify their limitations in the presence of stochastic noise and proposing improvements to enhance their applicability.
Keywords Policy Gradient Methods $\cdot$ Sample Complexity $\cdot$ Linear Optimal Control $\cdot$ Zeroth-order Optimization
1 Introduction
Reinforcement learning [1, 2, 3] has demonstrated a significant influence across a wide range of applications. A key concept within reinforcement learning is policy optimization, whereby the policy is parameterized and directly optimized over based on a predefined performance metric [4]. Several successful policy optimization methods have been developed, including policy gradient [5], actor-critic [6], proximal policy optimization [7]. This work focuses on policy gradient (PG) methods, which are based on the simple idea of minimizing a cost function over the parameterized policy by improving performance through a gradient descent-type update. Studying the convergence properties of PG methods, particularly their global convergence to the optimal policy, is an active area of research [8, 9, 10, 11, 12, 13, 14, 15]. For instance, PG methods have been applied in [11] to solve $\mathcal{H}_{2}$ cost function subject to $\mathcal{H}_{\infty}$ constraints, and they have been applied in [12] to Markov decision processes.
The application of reinforcement learning to the linear quadratic regulator (LQR) problem has been extensively explored due to its value as an analytically tractable benchmark, making it ideal for studying reinforcement learning in environments with continuous state and action spaces [16, 17, 9, 18, 19, 20, 21, 22]. As a consequence, PG methods have gained attention in recent studies focusing on the LQR problem. Research on these methods increased significantly after the work in [9] demonstrated global convergence of PG methods applied to the deterministic LQR problem. The work [13] explores a primal-dual policy gradient method to solve the constrained LQR problem. In [14], the performance limitations of PG applied to LQR problem are studied from a control-theoretic perspective.
Solving the LQR problem requires the availability of a model of the system, as it is assumed in [11, 12, 13, 14]. However, in real-world applications, a complete system description is often unavailable, making it necessary to combine policy optimization schemes with data-driven techniques. To address this challenge, various approaches have been developed. For instance, methods like those in [23, 24, 25] combine system identification with model-based LQR design and use regret as a performance metric of the learning process. It is also possible to use the same two-step approach to solve the LQR problem with gradient-based schemes. The work [26] integrates recursive least squares with policy iteration optimization, while [27] combines recursive least squares with policy gradient methods. Both works [26, 27] assume noise-free data, whereas [28, 29] extend these frameworks to bounded-noise settings. These approaches are termed indirect data-driven methods, because first data is used to estimate a model, which is then integrated into a model-based certainty equivalence design.
Direct data-driven methods instead use data directly to determine the policy, by-passing the intermediate model estimation process [30]. In [9], model-free PG methods are proposed using zeroth-order optimization to estimate gradients for the LQR cost from finite but noise-free system trajectories. The work [31] proposed a novel gradient estimation method for noise-free systems. The work [32] uses PG methods to solve the continuous-time state feedback LQR problem. Similarly, in [33], PG methods are applied to a continuous-time output feedback system using zeroth-order gradient estimation alongside variance reduction techniques. The work [34] focuses on stabilizing the unknown system using policy gradient methods. In the aforementioned direct data-driven works [9, 31, 32, 33, 34], only the case where the system’s trajectories are noise-free is studied. However, it is essential to analyze what happens when gradients and other objects involved in the policy update are estimated using noisy data. In [10], model-free PG methods are applied to linear systems with multiplicative noise, addressing uncertainty in the system matrix. The works [10, 33, 34] all build on [9], leveraging the inherent robustness of PG methods: informally, when the gradient estimation error is sufficiently small, one can still guarantee a decrease in the cost function. However, in all those works [9, 10, 33, 34], stochastic uncertainty is not considered, and the analysis assumes that the trajectory data is strictly bounded. Another line of research tackles the model-free LQR problem differently. The work [35] derives convergence properties by carefully selecting step sizes when facing uncertainty of the gradient. Follow-up works [36, 37] refine the analysis by incorporating alternative gradient estimation techniques, yielding improved sample efficiency and robustness guarantees. These results [35, 36, 37] assume stochastic but bounded noise. There are also studies considering stochastic unbounded noise. For example, [38] applies PG methods to finite-horizon LQR, while [39, 40] propose model-free first-order methods, which estimate individual components of the gradient using least squares and differ from the aforementioned zeroth-order approaches. To date, there is no comprehensive analysis that jointly addresses the uncertainty introduced by the estimation process using the zeroth-order method and its propagation through the algorithm, and provides probabilistic convergence guarantees and sample complexity bounds for the infinite-horizon LQR problem.
In this work, we analyze the sample complexity and robustness to noise of PG methods for the model-free LQR with additive stochastic noise. Due to the presence of noise with unbounded support, we consider an average infinite-horizon cost. We first characterize the properties of this cost function and analyze model-based versions of a few PG algorithms. For the policy gradient descent (PGD) and natural policy gradient (NPG) methods, we propose an adaptive stepsize scheme that has a provable faster convergence rate compared to the fixed step size methods discussed in [9, 10, 33]. Then, for both the model-free policy gradient descent and natural policy gradient methods, we employ zeroth-order optimization to approximate the gradient and covariance from noisy data. To handle the challenge of unbounded gradient estimates due to stochastic noise, we introduce an event-based analysis, extending the approach of [38], while relaxing the reliance on gradient upper bounds. We provide convergence guarantees for these model-free PG methods with a sample complexity that takes into account the noisy source of estimation error. Furthermore, we introduce and analyze a variance reduction technique within the zeroth-order optimization framework, and prove that this achieves a guaranteed improvement in the sample complexity. Even though variance reduction techniques for model-free LQR were already presented in [33], what we propose here is a new contribution compared to prior works, as we deal with stochastic data and we are able to provide a provable improvement in the number of required samples. Finally, we compare the guarantees of model-free PG methods with and without noisy data. This demonstrates the importance of considering noise, offers guidance on parameter tuning in applications, and also illustrates the improvement of introducing adaptive step sizes and variance reduction. Our analysis shows that model-free PG with stochastic data enjoys qualitatively similar convergence guarantees to those in the noise-free scenario, but it points out limitations in terms of the additional number of samples required for given accuracy, reduced step size ranges, and convergence rate.
The paper is organized as follows. Section 2 introduces the problem setting and provides some preliminaries. Section 3 studies convergence properties of a few representative PG methods for the average cost case and proposes their extension to adaptive stepsizes. Section 4 investigates convergence and sample complexity of model-free versions of gradient descent and natural policy gradient methods by placing particular emphasis on the effect of noisy data on the algorithms, together with adaptive step sizes and variance reduction techniques. Section 5 exemplifies the main findings of the work by observing the behavior of the analyzed PG methods under different noise levels and algorithmic options. Section 6 serves as a concluding summary of the work.
Notations
We denote by $A\succeq 0$ and $A\succ 0$ a positive semidefinite and positive definite matrix $A$, respectively. The symbol $\lambda_{1}(A)$ denotes the smallest eigenvalue of the matrix $A$. For matrices, $\lVert\cdot\rVert_{F}$ and $\lVert\cdot\rVert$ denote the Frobenius norm and induced $2$-norm, respectively. For vectors, $\lVert\cdot\rVert$ denotes the Euclidean norm. $I$ represents the identity matrix. $\mathbb{Z}_{+}$ is the set of non-negative integers. The operation $vec(A)=[a_{1}^{\top},a_{2}^{\top},...,a_{n}^{\top}]^{\top}$ stacks the columns of matrix $A$ into a vector. $\mathcal{N}(0,\Sigma)$ denotes a Gaussian distribution with 0 mean and covariance $\Sigma\succ 0$.
2 Preliminaries
In this work, we consider the following averaged infinite horizon optimal control problem, where the plant is subject to additive stochastic noise:
$\min_{u_{t}}$ $\lim_{T\rightarrow+\infty}\frac{1}{T}\mathop{\mathbb{E}}_{x_{0},w_{t}}\sum_{t=0}^{T-1}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}\right),$ (1a)
$\mathrm{s.\,t.}$~ $x_{t+1}=Ax_{t}+Bu_{t}+w_{t},w_{t}\sim\mathcal{N}(0,\Sigma_{w}),x_{0}\sim\mathcal{D},$ (1b)
where $x_{t}\in\mathbb{R}^{n_{x}}$ is the system state and $u_{t}\in\mathbb{R}^{n_{u}}$ is the system input; $A\in\mathbb{R}^{n_{x}\times n_{x}}$, $B\in\mathbb{R}^{n_{x}\times n_{u}}$, $(A,B)$ is unknown but stabilizable, which is a standard assumption in data-driven control [41, 42]; $Q,R\succ 0$ are the weighting matrix. We define the state covariance at the initial time as $\Sigma_{0}:=\mathop{\mathbb{E}}\limits_{x_{0}}\left[x_{0}x_{0}^{\top}\right]$ and then $\mathcal{D}:=\mathcal{N}(0,\Sigma_{0})$.
The input $u_{t}$ is parameterized as a linear state feedback control with gain $K$, i.e. $u_{t}=Kx_{t}$. The optimal control cost only depends on $K$ and is denoted by $C$:
$C(K):=\lim_{T\rightarrow+\infty}\frac{1}{T}\mathop{\mathbb{E}}_{x_{0},w_{t}}\sum_{t=0}^{T-1}x_{t}^{\top}\underbrace{\left(Q+K^{\top}RK\right)}_{=:Q_{K}}x_{t}.$ (2)
Remark 1 (Weighting matrices $Q,R$)
Prior knowledge of the weighting matrices $Q$ and $R$ is not strictly required for implementing the model-free policy gradient algorithms investigated here. While the analyses presented here assume their knowledge, it suffices to have access to the empirical cost, which can be estimated from the closed-loop trajectories of system (1b) used to estimate gradient information.
2.1 Properties of $C$
In this section, we introduce basic properties of problem (1) that are particularly relevant when policy gradient methods are used to solve it. While these results are mostly well-known, they form the foundation for the main results of this work in Sections 3 and 4. The state response $x_{t}$ for $t\geq 1$ associated with any gain $K$ is given by:
$x_{t}=(A_{K})^{t}x_{0}+\sum_{k=0}^{t-1}(A_{K})^{t-k}w_{k-1}.$ (3)
We define the set $\mathcal{S}$ as the set of matrices $K\in\mathbb{R}^{n_{x}\times n_{u}}$ that stabilizes the system $(A,B)$, meaning that the matrix $A_{K}:=A+BK$ is Schur stable:
$\mathcal{S}:=\left\{K\in\mathbb{R}^{n_{x}\times n_{u}}|\lVert A_{K}\rVert<1\right\}.$ (4)
Now, we introduce some important definitions associated with $K\in\mathcal{S}$. The covariance matrix at time $t$ is defined as $\Sigma_{t}:=\mathop{\mathbb{E}}\limits_{x_{0},w_{t}}[x_{t}x_{t}^{\top}]$. From this definition and the state response given in (3), we obtain:
$\Sigma_{t+1}=A_{K}\Sigma_{t}A_{K}^{\top}+\Sigma_{w},\quad t\in\mathbb{Z}_{+}.\\$ (5)
The average covariance matrix associated with $K\in\mathcal{S}$ is defined as:
$\Sigma_{K}:=\lim_{T\rightarrow+\infty}\frac{1}{T}\sum_{t=0}^{T-1}\Sigma_{t}.$ (6)
Then from (5) and (6), we have:
$\Sigma_{K}=\Sigma_{w}+A_{K}\Sigma_{K}A_{K}^{\top}=\sum_{t=0}^{\infty}A_{K}^{t}\Sigma_{w}A_{K}^{t~\top}.$ (7)
The cost function $C(K)$ associated with $K\in\mathcal{S}$ can be equivalently expressed as:
$C(K)=\mathrm{Tr}\left(P_{K}\Sigma_{w}\right)=\mathrm{Tr}\left(Q_{K}\Sigma_{K}\right),$ (8)
where $P_{K}$ is defined as the solution of the following Lyapunov equation:
$P_{K}=Q_{K}+A_{K}^{\top}P_{K}A_{K}=\sum_{t=0}^{\infty}A_{K}^{t~\top}Q_{K}A_{K}^{t}.$ (9)
From the expression of the cost function (8) and well-known results [43], the optimal $K^{*}$, which minimizes the cost function $C$, is given by:
$K^{*}$ $=-(R+B^{\top}P_{K^{*}}B)^{-1}B^{\top}P_{K^{*}}A,$ (10a)
$P_{K^{*}}$ $=Q+A^{\top}P_{K^{*}}A-A^{\top}P_{K^{*}}B(R+B^{\top}P_{K^{*}}B)^{-1}B^{\top}P_{K^{*}}A.$ (10b)
The average covariance matrix associated with the optimal $K^{*}$ is denoted as $\Sigma_{K^{*}}$. Using (8), the gradient of $C(K)$ with $K\in\mathcal{S}$ can be expressed as follows [14]:
$\nabla C(K)=2E_{K}\Sigma_{K},$ (11)
where $E_{K}:=\left(R+B^{\top}P_{K}B\right)K+B^{\top}P_{K}A$.
It is known that the cost function $C$ defined in (2) is generally non-convex [9]. In the context of non-convex optimization, the convergence to the optimal solution of gradient descent cannot usually be guaranteed. However, the cost function $C(K)$ satisfies a special property known as gradient domination.
Lemma 1 (Gradient Domination)
The function $C$ on the set $\mathcal{S}$ is gradient dominated. That is, for any $K\in\mathcal{S}$, the following inequality holds:
$C(K)-C(K^{*})\leq\mu\lVert\nabla C(K)\rVert_{F}^{2},$ (12)
with $\mu:=\frac{1}{4}\lVert\Sigma_{K^{*}}\rVert\lVert\Sigma_{w}^{-2}\rVert\lVert R^{-1}\rVert$.
From (12) and the fact that $\Sigma_{w}\succ 0$, it follows that the cost function $C$ has a unique minimizer and no other stationary points. The proof of Lemma 1 is provided in Appendix A.1, and is a straightforward extension of [9, Lemma 3] to the averaged infinite horizon setting considered here (2).
Lemma 2 (Almost Smoothness on $\mathcal{S}$)
For any $K,K^{\prime}\in\mathbb{S}$, the following inequality holds:
$\begin{split}\left\lvert C(K^{\prime})-C(K)-2\mathrm{Tr}\left(({K^{\prime}}-{K})^{\top}E_{K}\Sigma_{K^{\prime}}\right)\right\rvert\leq\lVert\Sigma_{K^{\prime}}\rVert\left\lVert R+B^{\top}P_{K}B\right\rVert\left\lVert K^{\prime}-{K}\right\rVert_{F}^{2},\end{split}$ (13)
The proof of Lemma 2 is provided in Appendix A.2, which is an extension of [9, Lemma 6]. To understand why (13) is referred to as almost smoothness, assume additionally that the term $\lVert\Sigma_{K}\rVert,\lVert R+B^{\top}P_{K}B\rVert$ can be upper bounded by a constant $L$, and that $K^{\prime}$ is sufficiently close to $K$ so that we can approximate $\Sigma_{K}$ with $\Sigma_{K^{\prime}}$, i.e., $\Sigma_{K}\approx\Sigma_{K^{\prime}}$. Then, recalling (13), we see that (13) is equivalent to:
$\left\lvert C(K^{\prime})-C(K)-\mathrm{Tr}\left(({K^{\prime}}-{K})^{\top}\nabla C(K)\right)\right\rvert\leq L\left\lVert K^{\prime}-{K}\right\rVert_{F}^{2},\quad\forall K,~K^{\prime}\in\mathcal{S.}$
which is the classic descent lemma for a smooth matrix function $C$. This approximation allows us to interpret the cost function as nearly smooth when the aforementioned assumptions are satisfied. Since we aim to leverage (13) to develop a gradient descent algorithm, it is crucial to quantify the relationship between $\Sigma_{K}$ and $\Sigma_{K^{\prime}}$ and find the upper bound $L$, which will be discussed in the following section.
2.2 Perturbation of $\Sigma_{K},C,\nabla C$
In this subsection, we demonstrate that the average covariance matrix $\Sigma_{K}$, the cost function $C(K)$, and its gradient $\nabla C(K)$ are locally Lipschitz continuous with respect to the policy $K$. This property is crucial for the analysis of model-free policy optimization.
Lemma 3 ($\Sigma_{K}$ Perturbation)
Suppose $K^{\prime},K\in\mathcal{S}$ are such that:
$\lVert K-K^{\prime}\rVert\leq\frac{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}{8C(K)\lVert B\rVert}=:h(C(K)),$
it holds that:
$\lVert\Sigma_{K}-\Sigma_{K^{\prime}}\rVert\leq\underbrace{8\left(\frac{C(K)}{\lambda_{1}(Q)}\right)^{2}\frac{\lVert B\rVert}{\lambda_{1}(\Sigma_{w})}}_{=:h_{\Sigma}(C(K))}\lVert K-K^{\prime}\rVert.$
The proof of Lemma 3 is given in Appendix A.3, which follows the proof of [9, Lemma 16] and uses the definition of $\Sigma_{K}$.
Lemma 4 ($C$ Perturbation)
Suppose $K^{\prime},K\in\mathcal{S}$ are such that:
$\lVert K-K^{\prime}\rVert\leq\min\{h\left(C(K)\right),\lVert K\rVert\},$
it holds that:
$\lVert C(K^{\prime})-C(K)\rVert\leq h_{C}(C(K))\lVert K-K^{\prime}\rVert,$
where $h_{C}(C(K))$ is defined in (55) in the proof.
Lemma 5 ($\nabla C$ Perturbation)
Suppose $K^{\prime}$ is such that:
$\lVert K-K^{\prime}\rVert\leq\min\{h(C(K)),\lVert K\rVert\},$
then there exists a polynomial $h_{\nabla}(C(K))$ such that:
$\lVert\nabla C(K)-\nabla C(K^{\prime})\rVert\leq h_{\nabla}(C(K))\lVert K-K^{\prime}\rVert,$
where $h_{\nabla}(C(K))$ is defined in (59) in the proof.
In summary, Lemma 3, Lemma 4 and Lemma 5 establish that $\Sigma_{K}$, $C$ and $\nabla C$ are locally Lipschitz continuous. The proofs of Lemma 4 and Lemma 5 are provided in Appendix A.4 and A.5, respectively. These proofs follow the approach used in [9, Lemma 24, Lemma 25] by adapting the expressions of $C$ and $\nabla C$.
3 Convergence of Model-based Policy Gradient
In this section, we show the global convergence properties of various model-based policy gradient methods, that is, algorithms that operate with perfect knowledge of the system matrices $(A,B)$. We consider three variants of policy gradient algorithms:
  • policy gradient descent: $K_{i+1}=K_{i}-\eta\nabla C(K_{i})$;
  • Natural Policy Gradient: $K_{i+1}=K_{i}-\eta\nabla C(K_{i})\Sigma_{K_{i}}^{-1}$;
  • Gauss-Newton: $K_{i+1}=K_{i}-\eta(R+B^{\top}P_{K_{i}}B)^{-1}\nabla C(K_{i})\Sigma_{K_{i}}^{-1}$,
where $i$ is the iteration index and $i\in\mathbb{Z}_{+}$. The Gauss-Newton method, when implemented with a step size $\eta=\frac{1}{2}$, becomes equivalent to the policy iteration algorithm, which is well-known for its quadratic convergence rate [44]. To implement the Gauss-Newton method, three terms, $(R+B^{\top}P_{K_{i}}B),\nabla C(K_{i}),\Sigma_{K_{i}}$ are necessary. In contrast, policy gradient descent only requires the gradient $\nabla C(K_{i})$, but exhibits only a linear convergence rate. The natural policy gradient method strikes a balance between the two, requiring less information than the Gauss-Newton method while providing faster convergence properties than the policy gradient descent method.
3.1 Policy gradient descent
In this subsection, we present the theorem guaranteeing the convergence of the model-based policy gradient descent method, based on the properties introduced in Sections 2.1 and 2.2.
Theorem 1 (Policy gradient descent with Adaptive Step Size)
Suppose the initial $K_{0}\in\mathcal{S}$, and consider the policy gradient descent iteration
$K_{i+1}=K_{i}-\eta_{i}\nabla C(K_{i}),\quad\forall i\in\mathbb{Z}_{+},$ (14)
where the step size satisfies $\eta_{i}\leq h_{\mathrm{PGD}}(C(K_{i}))$, and $h_{\mathrm{PGD}}(C(K_{i}))$ is defined in (62) in Appendix B.1. Then, the following relationship holds:
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}\right)(C(K_{i})-C(K^{*})).$ (15)
For $\eta_{i}=h_{\mathrm{PGD}}(C(K_{i})),i\in\mathbb{Z}_{+}$ and given any accuracy gap $\epsilon>0$, if the number of iterations $N$ satisfies:
$N\geq\frac{\lVert\Sigma_{K^{*}}\rVert}{2{\eta_{0}}\lambda_{1}(R)\lambda^{2}_{1}(\Sigma_{w})}\log\frac{C(K_{0})-C(K^{*})}{\epsilon},$
then
$C(K_{N})-C(K^{*})\leq\epsilon.$
The proof of Theorem 1 can be found in Appendix B.1. From Theorem 1, we see that the policy gradient descent method can achieve any desired accuracy $\epsilon$ within a finite number of iterations $N$ and converges linearly. According to Theorem 1, the step size is adaptive according to the cost $C(K_{i})$ at the current iterate $K_{i}$.
Remark 2 (Advantage of Adaptive Step Size over Fixed Step Size)
We can further characterize the advantage of using adaptive step sizes as proposed in Theorem 1. In [9], a fixed step size $\eta=h_{\mathrm{PGD}}(C(K_{0}))$ was proposed. While for both choices of step sizes, we can provide that $C(K_{i+1})\leq C(K_{i})$ from Theorem 1. We can also show that, given the same initial $K_{0}$ and together with the expression of $h_{\mathrm{PGD}}$, $\eta=\eta_{0}\leq\eta_{i},\forall i\in\mathbb{Z}_{+}$. This leads to the following inequality:
$0<\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}\right)\leq\left(1-\frac{2\eta_{0}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}\right)<1,\forall i\in\mathbb{Z}_{+}.$
This allows the improvement achieved with an adaptive step size to be quantified.
3.2 Natural Policy Gradient
The natural policy gradient method adjusts the standard policy gradient by considering the geometry of the parameter space through the Fisher information matrix [45]. The update rule for the natural policy gradient is given by:
$\theta\leftarrow\theta-\eta G_{\theta}^{-1}\nabla C(\theta),$ (16)
where $\theta$ is a vector of optimization variables and $G_{\theta}$ is the Fisher information matrix associated with $\theta$, and $\eta$ is the step size. An important special case is optimizing over a linear policy with additive Gaussian noise, expressed as:
$u_{t}=\pi_{K}(u_{t}|x_{t})=\mathcal{N}(Kx_{t},\alpha^{2}I),$ (17)
where $K\in\mathbb{R}^{n_{x}\times n_{u}}$ represents the optimization variables and $\alpha^{2}$ is the fixed noise variance, which can be set arbitrarily. Without loss of generality, we set $\alpha=1$. Since the optimization variable $\theta$ in (16) is a vector, and $K$ in (17) is a matrix, we define $\bar{K}$ as the vectorized form of $K$, such that $\theta=\bar{K}$:
$K\in\mathbb{R}^{n_{x}\times n_{u}}=\left[\begin{array}[]{cc}k_{1}\\ k_{2}\\ ...\\ k_{n_{x}}\end{array}\right]\rightarrow[k_{1},k_{2},...,k_{n_{x}}]^{\top}=:\bar{K}\in\mathbb{R}^{n_{x}n_{u}}.$
In this work, we define the Fisher information matrix $G_{\bar{K}}$ as:
$\hat{G}_{\bar{K}}:=\lim_{T\rightarrow+\infty}\frac{1}{T}\sum_{t=0}^{T}~\mathbb{E}\left[\nabla\mathop{log}\pi_{\bar{K}}(u_{t}|x_{t})\nabla\mathop{log}\pi_{\bar{K}}(u_{t}|x_{t})^{\top}\right].$ (18)
Given this definition, the natural policy gradient update, incorporating the Fisher information matrix $G_{\bar{K}}$, is given by:
$K_{i+1}=K_{i}-\eta\nabla C(K_{i})\Sigma_{K_{i}}^{-1}$ (19)
The explicit derivation of (19) starting from (16) and using the particular policy parametrization (17) is given in Appendix B.2. We now present the theorem that establishes the convergence guarantee of the natural policy gradient method
Theorem 2 (Natural Policy Gradient with Adaptive Step Size)
Suppose the initial $K_{0}\in\mathcal{S}$, and consider the natural policy gradient iteration
$K_{i+1}=K_{i}-\eta_{i}\nabla C(K_{i})\Sigma_{K_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$ (20)
where the step size satisfies
$\eta_{i}\leq\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}(\Sigma_{w})}}.$ (21)
Then the following relationship holds:
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}\right)(C(K_{i})-C(K^{*})).$ (22)
For
$\eta_{i}=\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}(\Sigma_{w})}},\quad\forall i\in\mathbb{Z}_{+},$ (23)
and given any accuracy gap $\epsilon>0$, if the number of iterations $N$ satisfies:
$N\geq\frac{\lVert\Sigma_{K^{*}}\rVert}{2\lambda_{1}(\Sigma_{w})}\left(\frac{\lVert R\rVert}{\lambda_{1}(R)}+\frac{\lVert B\rVert^{2}C(K_{0})}{\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}\right)\log\frac{C(K_{0})-C(K^{*})}{\epsilon},$
then
$C(K_{N})-C(K^{*})\leq\epsilon.$
The proof of Theorem 2 relies on Lemma 2 and follows the procedure outlined in [9, Lemma 15]. Theorem 2 establishes the global convergence properties of the natural policy gradient method, when an appropriate step size $\eta_{i}$ is selected. Similar to the observation in Remark 2, the introduction of adaptive step sizes improves the convergence rate, as the step sizes increase adaptively with the decreasing of the cost.
3.3 Gauss-Newton Method
We now present the theorem that provides the convergence guarantee for the Gauss-Newton method.
Theorem 3 (Gauss-Newton Method)
Suppose the initial $K_{0}\in\mathcal{S}$, and consider the Gauss-Newton iteration
$K_{i+1}=K_{i}-\eta(R+B^{\top}P_{K_{i}}B)^{-1}\nabla C(K_{i})\Sigma_{K_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$ (24)
with $\eta\leq\frac{1}{2}$. Then, the following relationship holds:
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{K}^{*}\rVert}\right)(C(K_{i})-C(K^{*})).$ (25)
For the maximum fixed step size $\eta=\frac{1}{2}$ and any accuracy gap $\epsilon>0$, if the number of iterations $N$ satisfies:
$N\geq\frac{\lVert\Sigma_{K}^{*}\rVert}{\lambda_{1}(\Sigma_{w})}\log\frac{C(K_{0})-C(K^{*})}{\epsilon},$
then
$C(K_{N})-C(K^{*})\leq\epsilon.$
The proof of Theorem 3 builds on the property shown in Lemma 2 and follows similar steps to the steps in [9, Lemma 14]. It is important to note that the choice of step size here is independent of the specific system parameters. To achieve the fastest convergence rate, one can always select the step size $\eta=\frac{1}{2}$, which is equivalent to the policy iteration algorithm.
The Gauss-Newton method requires knowledge of term $(R+B^{\top}P_{K}B)$, making it incompatible within the zeroth-order optimization frameworks in the model-free setting discussed in the following section. For implementation details of a model-free policy iteration method, please refer to [42], which employs least squares to estimate the necessary matrices from online data. However, it is important to note that [42] does not provide any convergence guarantees under noisy data conditions. Developing a sample-based estimator for the term $(R+B^{\top}P_{K}B)$ and establishing convergence guarantees for the Gauss-Newton method under noisy data conditions remains a key area for future research.
3.4 Lack of robustness of model-based PG methods
The guarantees on PG methods reviewed in the previous sections rely on the availability of an exact system model. When this is not available, gradients and covariances are estimated from data trajectories, and for plants such as (1b), they will be subject to stochastic errors. We exemplify with a simple toy example the performance of a policy GD method when noisy gradients are used. We consider a linear time-invariant (LTI) system described by the following dynamics [17, 42]:
$x_{t+1}=\underbrace{\left[\begin{array}[]{ccc}1.01&0.01&0\\ 0.01&1.01&0.01\\ 0&0.01&1.01\end{array}\right]}_{A}x_{t}+\underbrace{\left[\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}\right]}_{B}u_{t}+w_{t}.$ (26)
The weighting matrices $Q$ and $R$ are set to
$Q=0.001I_{3},\quad R=I_{3}$ (27)
The initial condition ${K}_{0}$ is chosen as the optimal gain for the LQR problem with $(A,B,50Q,R)$. To model gradient estimation errors due to noisy data, we define the gradient estimate as follows:
$\hat{\nabla}C(K):=\nabla C(K)+\Delta,$ (28)
where $\Delta\in\mathbb{R}^{3\times 3}$ is a matrix with entries $\Delta_{ij}$ sampled according to the distribution $\mathcal{N}(0,\sigma^{2})$, where the values of $\sigma\geq 0$ will be discussed later. The policy gradient descent algorithm updates the gain ${K}_{i}$ using (14) with fixed step size and the gradient estimate.
The simulation results are illustrated in Figure 1, where the y-axis represents the average obtained from a Monte Carlo simulation over 10,000 data samples.
Figure 1: Performance of Gradient Descent with inexact noisy gradient for different values of step size and noise
When the gradient is exactly known (i.e. $\sigma=0$, black dotted line), we can use theoretical bounds on the step size to guarantee convergence of the gradient descent algorithm. In this case, the converging behavior of $\eta=0.12$ to the optimal policy is shown as an example. When the gradient estimation is affected by a small amount of noise ($\sigma=0.03$, blue dashed line), the same step size yields convergence but to a suboptimal value of gain $K$. As the noise level increases ($\sigma=0.6$, green dash-dotted line), the algorithm starts showing non-converging behavior. In this case, reducing the step size addresses the issue, as shown by the red solid line where the step size is empirically decreased to $0.01$. It is worth observing that this reduction in step size results in a slower convergence rate.
These results exemplify that the convergence guarantees obtained in the model-based scenario no longer hold, when PG algorithms are run using finite noisy data trajectories to compute gradients (or equivalently covariances). Specifically, the algorithm might converge to suboptimal solutions or even diverge. It is also shown that, in the latter case, by properly modifying the algorithms’ parameters (such as the step size), converging behaviors can be achieved.
To characterize the fundamental limitations of such algorithms and improve their performance it is then paramount to: quantify the uncertainty on gradients and covariances due to the use of finite noisy data; and analyze how this uncertainty propagates in the policy optimization algorithm. To the best of the authors’ knowledge, there is no holistic analysis that captures all these aspects and provides guarantees for model-free PG algorithms in the presence of noisy trajectory data. The main technical contribution of this work is to bridge this gap by providing: probabilistic guarantees of convergence and suboptimality bounds; closed-form expressions for the algorithm’s parameters (e.g. step size, variance reduction) that guarantee converging behavior.
4 Convergence of Model-free Policy Gradient
In this section, we analyze two variants of model-free policy gradient methods: policy gradient descent and natural policy gradient. Unlike model-based approaches, the system dynamics $(A,B)$ are unknown in the model-free setting. Instead of relying on system identification, we use zeroth-order optimization, where gradients and covariance estimates are estimated directly from trajectory data. In Section 4.1, we begin by analyzing estimation algorithms for two fundamental components utilized in the proposed methods: the cost function gradient and the covariance. We derive novel finite-sample error bounds that explicitly account for the effects of noisy data. Building on these quantitative bounds, Section 4.2 introduces a variance reduction technique to provably reduce the required number of rollouts. Leveraging these results, we establish probabilistic convergence guarantees for PGD (Section 4.3) and NPG (Section 4.4), emphasizing the more general scenario with time-adaptive step sizes. In Section 4.3 and 4.4, we denote by $\hat{K}_{i}$ the gain at the $i$-th iteration, updated using gradient estimates.
4.1 Model-free Gradient and Covariance Estimation
In the model-free setting, zeroth-order optimization techniques [46, 47] can be utilized to approximate the gradient and average covariance matrix using only function evaluations. The procedures for gradient and covariance estimation are outlined in Algorithm 1.
Algorithm 1 Gradient and covariance estimation
0: Gain matrix $K\in\mathcal{S}$, number of rollouts $n$, rollout length $l$, exploration radius $r$, an upper bound of the initial state $L_{0}$
for k=1,,nk=1,...,n do
   1. Generate a sample gain matrix $\bar{K}_{k}=K+U_{k}$, where $U_{k}$ is drawn uniformly at random over matrices of compatible dimensions with Frobenius norm $r$;
   2. Generate a sample initial state $x_{0}^{(k)}$ with $\lVert x_{0}^{(k)}\rVert\leq L_{0}$;
   3. Excite the closed-loop system with $u_{t}^{(k)}=\bar{K}_{k}x_{t}^{(k)}$ for $l$-steps starting from $x_{0}^{(k)}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$ originating from (1b);
   4. Collect the empirical cost estimate $\hat{C}_{\bar{K}_{k}}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~\top}(Q+\bar{K}^{\top}_{k}R\bar{K}_{k})x_{t}^{(k)}$ and the empirical covariance matrix $\hat{\Sigma}_{k}=\frac{1}{l}\sum_{t=0}^{l-1}~x_{t}^{(k)}x_{t}^{(k)~\top}$;
end for
return  Gradient estimate $\hat{\nabla}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{\bar{K}_{k}}U_{k}$ and covariance estimate $\hat{\Sigma}_{K}:=\frac{1}{n}\sum^{n}_{k=1}\hat{\Sigma}_{k}$.
The empirical cost $\hat{C}_{\bar{K}_{k}}$ can be computed either by using the weighting matrix or directly from performance measurements, without requiring knowledge of $Q$ and $R$, as stated in Remark 1. We present two theorems to analyze the estimation errors of both $\hat{\nabla}C(K)$ and $\hat{\Sigma}_{K}$, obtained from Algorithm 1. The estimation error of $\hat{\nabla}C(K)$, denoted by $\epsilon$, together with its failure probability $\delta$, arises from several sources:
  • $\epsilon_{d}$: error caused by the noisy data;
  • $\epsilon_{l}$: error due to the finite rollout length compared to the infinite horizon cost;
  • $\epsilon_{n}$: error from using a finite number of rollouts in zeroth-order optimization;
  • $\epsilon_{r}$: difference between the gradient of the original function and the smoothing function $C_{r}(K)=\mathbb{E}_{U\sim\mathbb{B}_{r}}\left[C(K+U)U\right]$.
The probability consists of the following three components:
  • $\delta_{n}$: concentration probabilities for gradient estimation of the smoothing function, linked to $\epsilon_{n}$;
  • $\delta_{d}$: concentration probabilities for the noisy data, linked to $\epsilon_{d}$;
  • $\delta_{x}$: probability that the state is not bounded.
Having identified the different sources contributing to the estimation error and its associated failure probability, we are now in a position to state the main theorem.
Theorem 4 (Error Bound of $\hat{\nabla}C(K)$)
Given an arbitrary tolerance $\epsilon>0$, which can be expressed as $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$, and an arbitrary probability $\delta\in(0,1)$, which can be expressed as $\delta=1-(1-\delta_{d})(1-\delta_{n})(1-\delta_{x})$ with $\delta_{x},\delta_{n},\delta_{d}\in(0,1)$, for a given $K\in\mathcal{S}$, the estimated gradient $\hat{\nabla}C(K)$ from Algorithm 1 enjoys the following bound:
$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq 1-\delta,$ (29)
if the parameters $r,l,n$ in Algorithm 1 satisfy:
$r\leq r_{\max}(C(K),\epsilon_{r}),$ $\quad l\geq l_{\min}(C(K),\epsilon_{l}),$ (30a)
$n\geq\max(N_{1}(C(K),\epsilon_{r},\epsilon_{l},\epsilon_{n},\delta_{n},\delta_{d}),$ $N_{2}(C(K),\epsilon_{r},\epsilon_{l},\epsilon_{n},\epsilon_{d},\delta_{x},\delta_{n},\delta_{d})),$ (30b)
where the detailed expressions of functions $r_{\max}$, $l_{\min}$, $N_{1}$ and $N_{2}$ are given in (92), (93), (94a), and (94b) in Appendix C.5, respectively.
The proof of Theorem 4 is given in Appendix C.5. Similar to the estimation error of the gradient, the estimation error of $\Sigma_{K}$,i.e., $\epsilon^{\prime}$ and its associated failure probability $\delta^{\prime}$ also originate from different sources, which we outline below: the error $\epsilon^{\prime}$ consists of three components:
  • $\epsilon^{\prime}_{l}$: error due to the finite rollout length compared to the infinite horizon average covariance;
  • $\epsilon^{\prime}_{n}$: error from noise in the data;
  • $\epsilon^{\prime}_{r}$: difference between average covariance associated with $K$ and $\bar{K}_{k}$.
The probability $\delta$ consists of two components:
  • $\delta^{\prime}_{n}$: concentration probabilities for the noisy data;
  • $\delta^{\prime}_{x}$: probability that the state remains bounded.
Theorem 5 (Error bound of $\hat{\Sigma}_{K}$)
Given an arbitrary tolerance $\epsilon^{\prime}>0$, which can be expressed as $\epsilon^{\prime}=\epsilon^{\prime}_{l}+\epsilon^{\prime}_{n}+\epsilon^{\prime}_{r}$, and an arbitrary probability $\delta^{\prime}\in(0,1)$, which can be expressed as $\delta^{\prime}=1-(1-\delta^{\prime}_{n})(1-\delta^{\prime}_{x})$ with $\delta^{\prime}_{n},~\delta^{\prime}_{x}\in(0,1)$, for a given $K\in\mathcal{S}$, the estimated average covariance $\hat{\Sigma}_{K}$ from Algorithm 1 enjoys the following bound:
$\mathbb{P}\left\{\lVert\hat{\Sigma}_{K}-\Sigma_{K}\rVert\leq\epsilon^{\prime}\right\}\geq 1-\delta^{\prime},$ (31)
if the parameters $r,l,n$ in Algorithm 1 satisfy:
$r\leq r^{\prime}_{\max}(C(K),\epsilon^{\prime}_{r}),l\geq l^{\prime}_{\min}(C(K),\epsilon^{\prime}_{l}),n\geq n^{\prime}_{\min}(C(K),\epsilon^{\prime}_{r},\epsilon^{\prime}_{l},\epsilon^{\prime}_{n},\delta^{\prime}_{x},\delta^{\prime}_{n}),$ (32)
where the detailed expressions of functions $r^{\prime}_{\max}$, $l^{\prime}_{\min}$ and $n^{\prime}_{\min}$ are given in (98) (99) and (100) in the Proof C.6, respectively.
The proof of Theorem 5 is given in Appendix C.6.
We analyze the effect of estimation errors caused by stochastic noise acting on (1b). This analysis requires the introduction of new tools and a significant extension of existing results, where previous works [9, 10, 33] assumed that the trajectory has bounded support. Moreover, while [9] assumes that all error terms are equal, i.e., $\epsilon_{r}=\epsilon_{n}=\epsilon_{l}$ and $\epsilon_{r}^{\prime}=\epsilon_{n}^{\prime}=\epsilon_{l}^{\prime}$, we present a more structured approach to the error decomposition. By selecting different values for $l$, $r$, and $n$ (or $l^{\prime}$, $r^{\prime}$, $n^{\prime}$), the same accuracy $\epsilon$ (or $\epsilon^{\prime}$) can be achieved, but with a more refined breakdown of the contributing error terms.
Based on the results of Theorem 4 and Theorem 5, we conclude that, with appropriate choices of $l$, $r$, and $n$, the estimates $\hat{\Sigma}_{K}$ and $\hat{\nabla}C(K)$ from Algorithm 1 can achieve any desired accuracy with a corresponding desired probability. These estimates then replace their model-based counterparts in model-free policy gradient methods. The error bounds established in this section serve as a foundation for the convergence guarantees of both model-free policy gradient descent and natural policy gradient methods, which are discussed in the next sections.
4.2 Model-free Gradient Estimation with Variance Reduction
In the previous section, we employed zeroth-order optimization to estimate the gradient of the cost at the current value of the policy. However, this approach often suffers from high variance, resulting in a slow learning process [48]. Using a baseline is a common variance reduction technique in policy gradient methods [49, 50]. In this section, we propose employing the finite-horizon cost function as a baseline and show its performance improvement. For a state-dependent baseline function $b(x)$, the estimated gradient with variance reduction $\hat{\nabla}C_{v}(K)$ is expressed as
$\hat{\nabla}C_{v}(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}(\hat{C}_{\bar{K}_{k}}-b(x_{0}^{(k)}))U_{k},$ (33)
where $U_{k},x_{0}^{(k)},\hat{C}_{\bar{K}_{k}},r$ are the same as those defined in Algorithm 1. The estimated gradient $\hat{\nabla}C_{v}(K)$ satisfies:
$\begin{split}\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}C_{v}(K)]=\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}C(K)]-\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\mathbb{E}_{x_{0}^{(k)},U_{k}}[b(x_{0}^{(k)})U_{k}].\end{split}$ (34)
Since $U_{k}$ and $x_{0}^{(k)}$ are independent (as described in Algorithm 1), then
$\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}C_{v}(K)]=\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}C(K)].$
The trace of the covariance matrix of vectorized matrix $vec(Z_{k}^{V})$ is given as:
$\mathrm{Tr}(\mathrm{Var}(vec(Z_{k}^{V})))=\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[\left\lVert Z_{k}^{V}\right\rVert_{F}^{2}\right]-\left\lVert\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[Z_{k}^{V}\right]\right\rVert_{F}^{2}.$ (35)
Next, we determine the optimal function $b^{*}$, which minimizes the trace of the variance matrix above:
$\mathrm{Tr}(\mathrm{Var}(vec(Z_{k}^{V})))=\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[\left\lVert(\hat{C}_{\bar{K}_{k}}-b(x_{0}^{(k)}))U_{k}\right\rVert_{F}^{2}\right]-\left\lVert\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[(\hat{C}_{\bar{K}_{k}}-b(x_{0}^{(k)}))U_{k}\right]\right\rVert_{F}^{2}.$ (36)
For any state-dependent baseline function $b$, since $U_{k}$ is i.i.d and zero mean, we have:
$\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[b(x_{0}^{(k)})U_{k}\right]=0.$
Consequently, the second term in (36) becomes independent of the choice of $b$, leaving only the first term in (36) to be minimized. To minimize the first term, the baseline function should be chosen to reduce the mean square error of $\hat{C}_{\bar{K}_{k}}$. Thus, the optimal baseline $b^{*}$ is given by:
$b^{*}(x_{0})=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{w_{t},U_{k}}\left[x_{t}^{~\top}(Q+({K}+U_{k})^{\top}R({K}+U_{k}))x_{t}\right].$ (37)
Since $U_{k}$, which was also already defined in Algorithm 1, enters nonlinearly in (37), it is hard to compute the expectation and thus to obtain a closed-form expression for $b^{*}(x_{0})$. While it is possible to estimate it directly from data, the lack of a closed-form solution makes it challenging to establish explicit performance improvement guarantees. Instead, a suboptimal solution $b_{s}$ is used:
$b_{s}(x_{0})=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{w_{t}}\left[x_{t}^{~\top}(Q+{K}^{\top}R{K})x_{t}\right],$ (38)
The rationale for this choice is to simplify the computation of the expected value in (37) so that the performance improvement can be provided. In the next part, we will demonstrate (Theorem 6) that using $b_{s}$ leads to a provable reduction in the number of rollouts required, compared to the standard gradient computation discussed in the previous section.
Algorithm 2 summarizes the procedure used to compute the baseline function from data.
Algorithm 2 Baseline function estimation $b_{s}(x_{0})$ for variance reduction
0: Gain matrix $K\in\mathcal{S}$, number of rollouts to estimate the baseline function $n_{v}$, rollout length $l$, initial state $x_{0}$
for $k=1,...,n_{v}$ do
   1. Excite the closed-loop system with $u_{t}=Kx_{t}$ for $l$-steps starting from $x_{0}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$;
   2. Collect the empirical finite-horizon cost estimate $\hat{C}^{V}_{k}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~\top}(Q+{K}^{\top}R{K})x_{t}^{(k)}$;
end for
return  Baseline function estimate $\hat{b}_{s}(x_{0}):=\frac{1}{n_{v}}\sum^{n_{v}}_{k=1}\hat{C}^{V}_{k}$.
Using the estimated baseline function $\hat{b}_{s}$, the gradient estimation algorithm can be formulated as follows:
Algorithm 3 Gradient estimation with variance reduction
0: Gain matrix $K\in\mathcal{S}$, number of rollouts $n_{b}$, rollout length $l$, exploration radius $r$, an upper bound of the initial state $L_{0}$
for $k=1,...,n_{b}$ do
   1. Generate a sample initial state $x_{0}^{(k)}$ with $\lVert x_{0}^{(k)}\rVert\leq L_{0}$;
  2. Estimate the baseline function $\hat{b}_{s}(x_{0}^{(k)})$ using Algorithm 2
   3. Generate a sample gain matrix $\bar{K}_{k}=K+U_{k}$, where $U_{k}$ is drawn uniformly at random over matrices of compatible dimensions with Frobenius norm $r$;
   4. Excite the closed-loop system with $u_{t}^{(k)}=\bar{K}_{k}x_{t}^{(k)}$ for $l$-steps starting from $x_{0}^{(k)}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$;
   5. Collect the empirical finite-horizon cost estimate $\hat{C}_{\bar{K}_{k}}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~\top}(Q+\bar{K}^{\top}_{k}R\bar{K}_{k})x_{t}^{(k)}$
end for
return  Gradient estimate $\hat{\nabla}C(K):=\frac{1}{n_{b}}\sum^{n_{b}}_{k=1}\frac{n_{x}n_{u}}{r^{2}}(\hat{C}_{\bar{K}_{k}}-\hat{b}_{s}(x_{0}^{(k)}))U_{k}$
Compared with Algorithm 1, Algorithm 3 introduces an additional step involving the baseline function to reduce the variance of the estimated gradient. The following theorem establishes that this modification leads to a provable improvement in the sample complexity required to achieve a given gradient accuracy with high probability. Here, the overall probability is expressed as two parts: $\tilde{\delta}_{x}$, representing the probability that the samples remain bounded, and $\tilde{\delta}_{v}$, capturing the concentration of the sample.
Theorem 6 (Performance Improvement)
Given the same tolerance $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$ and probability $\delta=1-(1-\delta_{d})(1-\delta_{n})(1-\delta_{x})$ introduced in Theorem 4, for a given $K\in\mathcal{S}$, then the estimated gradient $\hat{\nabla}C(K)$ from Algorithm 3 enjoys the following bound:
$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq 1-\delta,$ (39)
if $n_{b},l$ and $r$ in Algorithm 3 satisfy:
$r$ $\leq r_{\max}(C(K),\epsilon_{r}),l\geq l_{\min}(C(K),\epsilon_{l}),$ (40a)
$n_{b}$ $\geq\max(N_{1},N_{3}(\hat{b}_{s}(x_{0}^{(k)}))$ (40b)
where the detailed expression of $N_{3}$ is given in (105) in Appendix C.7 and $r_{\max}$, $l_{\min}$ and $N_{1}$ were introduced in Theorem 4. Moreover, let the desired probability $\delta_{v}\in(0,1)$ be expressed as $\delta_{v}=1-(1-\tilde{\delta}_{v})(1-\tilde{\delta}_{x})$. If the number of rollouts $n_{v}$ to estimate the baseline function (as defined in Algorithm 2) satisfies
$n_{v}\geq\tilde{n}_{\min}(C(K),l,\tilde{\delta}_{v},\tilde{\delta}_{x}),$ (41)
then:
$\mathbb{P}\left\{N_{2}\geq N_{3}\right\}\geq 1-\delta_{v},$ (42)
where $N_{2}$ was first introduced in (30b) in Theorem 4 and the detailed expression of $\tilde{n}_{\min}$ is provided in (110) in Appendix C.7.
The proof of Theorem 6 is given in Appendix C.7. Based on Theorem 6, we can compare the performance of Algorithm 1 without variance reduction and Algorithm 3 with variance reduction. The exploration radius $r_{\max}$ and rollout length $l_{\min}$ (40a) remain the same for both algorithms. The number of rollouts (40b) is determined by the maximum of two arguments. $N_{1}$ remains unchanged across the two algorithms, as it depends on the smoothing function (see (67))and the original function. From the second part of Theorem 6 ((42)), it then follows that the required number of rollouts is at least non-increasing when the baseline function estimates are sufficiently accurate. As noted in Theorem 4, $\delta_{x}$ represents the probability that the data samples remain bounded. Increasing $\delta_{x}$ leads to a larger bound for the data samples. This upper bound is utilized in the matrix concentration inequality to determine the required number of rollouts. From (94a) and (94b), we observe that, for large $\Sigma_{w}$ and thus large noise, the required number of rollouts is determined by $N_{2}$ (in Algorithm 1) and $N_{3}$ (in Algorithm 3). In other words, the proposed variance reduction gives a provable reduction of the number of rollouts (and thus of samples) for scenarios with large noise, which is intuitive and expected.
4.3 Policy gradient descent with Adaptive Step Size
Building on the gradient uncertainty quantified in the previous section, we now proceed to study the convergence of the model-free policy gradient descent algorithm.
Algorithm 4 Model-free policy gradient descent with adaptive step size
0: An initial stabilizing gain matrix $\hat{K}_{0}$, desired accuracy $\epsilon$ and probability $\delta$.
[…]
  1. Compute the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ based on Theorem 4 or 6 to achieve the accuracy stated in Theorem 7.
  2. Use Algorithm 1 or 3 to estimate the gradient $\hat{\nabla}C(\hat{K}_{i})$.
   3. Update the gradient as $\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i})$, with $\eta_{i}\leq h_{\mathrm{PGD}}(C(\hat{K}_{i}))$.
end for
Theorem 7 (Model-free policy gradient descent with Adaptive Step Size)
Suppose the initial $\hat{K}_{0}\in\mathcal{S}$, and consider the policy gradient descent with adaptive step size:
$\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i}),\quad\forall i\in\mathbb{Z}_{+},$ (43)
where $\hat{\nabla}C(\hat{K}_{i})$ is the gradient estimate from Algorithm 1 and $0<\eta_{i}\leq h_{\mathrm{PGD}}(C(\hat{K}_{i}))$, where the function $h_{\mathrm{PGD}}$ was introduced in Theorem 1. Given any accuracy $\epsilon>0$, and $\sigma\in(0,1)$, define ${\eta_{\mathrm{PGD}}}:=\inf\limits_{i}\eta_{i}$ and the number of iterations $n_{\mathrm{PGD}}$:
$n_{\mathrm{PGD}}=\frac{\lVert\Sigma_{K^{*}}\rVert}{2(1-\sigma)\eta_{\mathrm{PGD}}\lambda_{1}(R)\lambda^{2}_{1}(\Sigma_{w})}\log\frac{C(\hat{K}_{0})-C(K^{*})}{\epsilon}.$
Given any probability $\delta\in(0,1)$ satisfying $\delta n_{\mathrm{PGD}}\in(0,1)$, assume the estimation error of the gradient $\hat{\nabla}C(\hat{K}_{i})$ satisfies:
$\mathbb{P}\left\{\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{h_{C}(C(\hat{K}_{i}))\lVert\Sigma_{K^{*}}\rVert}\right\}\geq 1-\delta,$
where $h_{C}$ was introduced in Lemma 4. Then for any $C(\hat{K}_{i})\geq C(K^{*})+\epsilon$, the following inequality holds:
$\mathbb{P}\left\{C(\hat{K}_{i+1})-C(K^{*})\leq\gamma_{i}(C(\hat{K}_{i})-C(K^{*}))\Big{|}C(\hat{K}_{i})\right\}\geq 1-\delta,$
where $\gamma_{i}:=1-(1-\sigma)\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}$ and $\gamma_{i}<1,\forall i\in\mathbb{Z}_{+}$.
As a result, the policy gradient descent method enjoys the following performance bound:
$\mathbb{P}\left\{\min_{i\in[0,n_{\mathrm{PGD}}]}C(\hat{K}_{i})-C(K^{*})\leq\epsilon\right\}\geq 1-\delta n_{\mathrm{PGD}}.$
To achieve the desired accuracy of the estimates stated in Theorem 7, the corresponding values of $l_{i},r_{i},n_{i}$, as defined in Algorithm 4, should be selected based on Theorem 4. Theorem 7 ensures that the cost will converge to the optimal solution with a predefined accuracy gap $\epsilon$, meaning $C(K^{*})+\epsilon$ will be reached with a certain probability. However, due to the estimation error in the gradient, further improvements in cost stop when the realized cost $C(\hat{K}_{i})\leq C(K^{*})+\epsilon$. To achieve higher accuracy, it is necessary to minimize the gradient estimation error. Nevertheless, the algorithm can only converge to the optimal when the gradient estimation is exact, which is not possible with a finite number of rollouts. The interpretation of the $\sigma$ is given in Remark 3. Regarding the step size $\eta_{i}$ stated in Theorem 7, from the expression of $h_{\mathrm{PGD}}$ given in (62), we observe that as the noise level $\Sigma_{w}$ increases, the step size $\eta_{i}$ decreases, reflecting the observation analyzed in Section 3.4.
In Algorithm 4, the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ are determined adaptively to the cost $C(\hat{K}_{i})$, i.e., they are computed online inside the for-loop. This design is aligned with Theorem 7, which shows that the cost function $C(\hat{K}_{i})$ decreases probabilistically with each iteration. As the cost $C(\hat{K}_{i})$ decreases, this leads to a reduction in the required rollouts $n_{i}$ and rollout length $l_{i}$, while the exploration radius $r_{i}$ increases. Alternatively, the algorithm can be modified to compute these parameters offline, where for all iterations $i$, $n_{i},l_{i},r_{i}$ are calculated once based on the initial cost $C(\hat{K}_{0})$.
4.4 Natural Policy Gradient with Adaptive Step Size
Analogous to the model-free policy gradient descent method, we now turn to analyzing the convergence of the natural policy gradient method.
Algorithm 5 Model-free natural policy gradient with adaptive step size
0: An initial stabilizing gain matrix $\hat{K}_{0}$, desired accuracy $\epsilon$ and probability $\delta$.
[…]
  1. Compute the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ based on Theorem 4 to achieve the accuracy stated in Theorem 8.
  2. Use Algorithm 1 to estimate the gradient $\hat{\nabla}C(\hat{K}_{i})$ and covariance $\hat{\Sigma}_{\hat{K}_{i}}$.
   3. Update the gradient as $\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}^{-1}_{\hat{K}_{i}}$ with $\eta_{i}\leq\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert C(\hat{K}_{i})}{\lambda_{1}(\Sigma_{w})}}$.
end for
Theorem 8 (Model-free Natural Policy Gradient with Adaptive Step Size)
Suppose the initial $\hat{K}_{0}\in\mathcal{S}$, and consider natural policy gradient with adaptive step size:
$\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$ (44)
where $\hat{\nabla}C(\hat{K}_{i})$ and $\hat{\Sigma}_{\hat{K}_{i}}$ are the gradient and covariance estimates from Algorithm 1 and $0<\eta_{i}\leq\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert C(\hat{K}_{i})}{\lambda_{1}(\Sigma_{w})}}$.
Given any accuracy $\epsilon>0$ and $\sigma\in(0,1)$, defining ${\eta_{\mathrm{NPG}}}:=\inf\limits_{i}\eta_{i}$ and the number of iterations $n_{\mathrm{NPG}}$:
$n_{\mathrm{NPG}}\geq\frac{\lVert\Sigma_{K^{*}}\rVert}{2(1-\sigma)\eta_{\mathrm{NPG}}\lambda_{1}(\Sigma_{w})}\log\frac{C(K_{0})-C(K^{*})}{\epsilon}.$
Given any probability $\delta\in(0,1)$ satisfying $\delta n_{\mathrm{NPG}}\in(0,1)$, assume that the estimation error of the gradient $\hat{\nabla}C(\hat{K}_{i})$ and the covariance $\hat{\Sigma}_{\hat{K}_{i}}$ satisfy:
$\mathbb{P}$ $\left\{\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}\right\}\geq\sqrt{1-\delta},$ (45a)
$\mathbb{P}$ $\left\{\lVert\hat{\Sigma}_{\hat{K}_{i}}-{\Sigma}_{\hat{K}_{i}}\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\sqrt{b_{\nabla}(C(\hat{K}_{i}))}}\right\}\geq\sqrt{1-\delta},$ (45b)
where $h_{\nabla}$ was introduced in Lemma 5 and $b_{\nabla}$ was defined in (56) in Appendix A.5. Then for any $C(\hat{K}_{i})\geq C(K^{*})+\epsilon$, the following inequality holds:
$\mathbb{P}\left\{C(\hat{K}_{i+1})-C(K^{*})\leq\kappa_{i}(C(\hat{K}_{i})-C(K^{*}))\Big{|}C(\hat{K}_{i})\right\}\geq 1-\delta,$
where $\kappa_{i}:=\left(1-(1-\sigma)\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{K}^{*}\rVert}\right)$ and $\kappa_{i}<1,\forall i\in\mathbb{Z}_{+}$.
As a result, the natural policy gradient method enjoys the following performance bound:
$\mathbb{P}\left\{\min_{i\in[0,n_{\mathrm{NPG}}]}C(\hat{K}_{i})-C(K^{*})\leq\epsilon\right\}\geq 1-\delta{n_{\mathrm{NPG}}},$
From Theorem 4, we can select the values of $l^{\nabla}_{i},r^{\nabla}_{i},n^{\nabla}_{i}$ to satisfy the requirement for the gradient estimation error, as indicated in (45a). Similarly, we can select $l^{\Sigma}_{i},r^{\Sigma}_{i},n^{\Sigma}_{i}$ to satisfy the requirement for the covariance estimation error, as described in (45b). To ensure both requirements are met, we can then set: $n_{i}\geq\max\{n^{\nabla}_{i},n^{\Sigma}_{i}\},l_{i}\geq\max\{l^{\nabla}_{i},l^{\Sigma}_{i}\},r_{i}\leq\min\{r^{\nabla}_{i},r^{\Sigma}_{i}\}$. These choices guarantee that the estimates of both the gradient and the covariance matrix satisfy their respective accuracy requirements. This combined selection of parameters is implemented in Algorithm 5 to ensure convergence with the desired accuracy and probability.
Remark 3 (Qualitative effect of gradient and covariance errors on convergence)
Model-free algorithms are extensions of model-based policy gradient methods. However, in the model-free setting, the convergence rate is typically slower, reduced by a factor of ($1-\sigma$) compared to the model-based case. This reduction is due to estimation errors in the gradient and covariance matrices, which are inherent in model-free methods as they rely on data-driven estimates rather than exact model information. When the estimated gradient closely matches the true gradient, the model-free policy gradient approaches the same convergence rate as the model-based policy gradient, corresponding to $\sigma\to 0$. However, if the estimation error becomes large, such that $\sigma\to 1$, further policy improvements become impossible, and the algorithm can no longer guarantee convergence to the optimal with the desired accuracy. This effect of gradient estimation error on the convergence rate and suboptimality gap aligns with our observations in Section 3.4, as illustrated by the black dotted line, blue dashed line, and green dash-dotted line in Figure 1.
4.5 Comparison of Model-free PG for Noise and Noise-free Case
In Algorithm 1, three key quantities, i.e. exploration radius $r$, rollout length $l$, and the required rollouts $n$, play a central role in determining the estimation errors on gradient and covariance estimation. The presence of stochastic noise acting on the system (1b) during data collection influences these quantities in the following ways:
[…]
  • Rollout length: The rollout length $l$ is different for the noise-free case and the noise case. In both scenarios, the rollout length plays a crucial role in estimating the gradient and covariance, whose true values are defined over an infinite horizon. However, finite approximations must be used. In the noise-free case, as described in [9, Lemma 23], the rollout length is determined by the desired accuracy, model parameters, and initial covariance $\Sigma_{0}$. In the presence of noise, as discussed in Lemma 6, the rollout length is influenced not only by $\Sigma_{0}$, but also by the noise covariance $\Sigma_{w}$. Therefore, the rollout length in the noisy case must account for both the system’s initial conditions and noise disturbances. These relationships are quantitatively explored in Lemma 6.
  • Required rollouts: The most significant difference between the noisy and noise-free cases comes from using the matrix concentration inequality for the trajectories consisting of noisy data. In the noise-free case, the boundedness of the data is guaranteed by an upper bound on the initial state, i.e., $\lVert x_{0}^{(k)}\rVert\leq L_{0}$ (see [9, Lemmas 27, 29]). In contrast, in the noisy case, particularly with unbounded Gaussian noise, the state remains bounded only with a certain probability (Appendix C.4). This boundedness in probability is then used to apply matrix concentration inequalities, providing an upper bound on the covariance of the samples(Appendix C.5 and [51, Appendix C.6]).
5 Simulation Results
In this subsection, we use numerical simulations  [^1] The Matlab codes used to generate these results (in Section 3.4 and Section 5) are accessible from the repository: https://github.com/col-tasas/2025-PGforLQRwithStochastic to demonstrate the benefits of incorporating the variance reduction technique and adaptive step sizes in model-free policy gradient algorithms and the overall effect of noise on them. The matrices $(A,B,Q,R)$ and $\hat{K}_{0}$ are the same as described in Section 3.4. The rollout length, exploration radius, and number of rollouts are set to l=100,r=0.04,n=1000l=100,~r=0.04,~n=1000, respectively. The simulation results are obtained from a Monte Carlo simulation over $5$ data samples.
5.1 Noise Level and Stepsize
The policy is updated using the policy gradient descent method in Algorithm 4. Figure 2 illustrates the relative suboptimality gap of the cost associated with the iterated policy $\hat{K}_{i}$ as a function of the iteration index $i$, for different noise levels and step sizes.
Figure 2: Effect of noise level on step size
From Figure 2, when the noise level is low, $\Sigma_{w}=10^{-4}I$, the step size is empirically chosen as $40$, as shown by the red dotted line. In this case, we observe a persistent suboptimality gap, which originates from the estimation error of the gradient inherent to the zeroth-order method from noisy data. When the noise level increases to $\Sigma_{w}=10^{-2}I$, a large step size ($\eta=6$) leads to divergence of the cost function (as shown in black dashed line). As stated in Theorem 7, when the noise level is higher, it is necessary to decrease the step size to ensure convergence. By reducing the step size to $\eta=0.3$, as illustrated by the blue solid line in the figure, the cost converges again with a suboptimality gap.
5.2 Effect of variance reduction on PGD
The step size $\eta$ is chosen based on the noise level: for $\Sigma_{w}=10^{-4}I$ it is set to $40$, while for $\Sigma_{w}=10^{-2}I$, it is reduced to $0.3$, qualitatively in accordance with the bounds established in Theorem 7. For the variance reduction technique, the number of rollouts to estimate the baseline function, $n_{b}$, is set to $200$.
Figure 3: PGD with/out variance reduction
Figure 3 illustrates the relative suboptimality gap of the cost associated with the iterated policy $\hat{K}_{i}$ as a function of the iteration index $i$. We compare two groups of curves corresponding to different noise levels: the blue dashed and black solid lines for the noise level $\Sigma_{w}=10^{-4}I$, and the green dash-dotted and magenta dotted lines for $\Sigma_{w}=10^{-2}I$. The blue dashed and green dash-dotted lines represent results without variance reduction. It can be observed that the performance improvement achieved through variance reduction is more pronounced in high-noise scenarios, which is consistent with the analysis in Theorem 6.
5.3 Model-free NPG with Adaptive Step Sizes
The policy is updated using the Natural policy gradient method in Algorithm 5. For the case of fixed step size it is chosen $\eta=\frac{a}{b+c\mathrm{Tr(P_{\hat{K}_{0}})}}$ which follows the structure of the step size derived in Theorem 8 and is applied uniformly across all noise levels $\Sigma_{w}={10^{-4}I,10^{-2}I,10^{0}I}$. The parameters $a$, $b$, and $c$ are selected empirically as $0.09$, 11, and $2$, respectively. For the case of adaptive step sizes, we take $\eta_{i}=\frac{a}{b+c\mathrm{Tr(P_{\hat{K}_{i}})}}$, where $\mathrm{Tr}(P_{K_{i}})$ reflects the current cost $C(\hat{K}_{i})$ associated with the policy at iteration $i$.
Figure 4: NPG with/out adaptive step size
Comparing the black solid line and the green dashed line in Figure 4, we observe that using adaptive step sizes yields a faster convergence rate than fixed step sizes, further highlighting the benefits of adaptive strategies. As shown in Figure 4 (red dotted, green dashed, and blue dot–dashed lines), the model-free NPG algorithm achieves nearly the same accuracy even as the noise $\Sigma_{w}$ increases, provided that $l,r,$ and $n$ remain unchanged. At higher noise levels, the estimation errors of both $\hat{\nabla}C(\hat{K}_{i})$ and $\hat{\Sigma}_{\hat{K}_{i}}$ grow. These errors are explicitly accounted for in our analysis (Theorems 4 and 5). Due to the natural policy gradient update rule (see (44)), the estimated gradient is multiplied by the inverse of the estimated covariance, which causes the errors in these two quantities to partially offset each other. If the analysis were performed directly on the product $\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}$ rather than treating them separately (as in Proof of Theorem 2), substantially stronger theoretical guarantees could be established. This observation points to an important direction for future research.
6 Conclusions
In this work, we applied the policy gradient method to the Linear Quadratic Regulator problem with stochastic noise, analyzing both model-based and model-free approaches. Convergence guarantees were established for the model-based gradient descent and natural policy gradient methods with adaptive step sizes, demonstrating their ability to converge to the global optimum. For the model-free gradient descent and natural policy gradient methods, we showed that they can converge to the optimal solution within any desired accuracy. However, it is crucial that the algorithm’s parameters are tuned based on factors such as the noise magnitude affecting the data trajectories. To improve convergence rates and reduce sample complexity, in addition to adaptive step sizes, we introduced a variance reduction technique. We also provided a qualitative discussion on the impact of noise on model-free policy gradient methods, offering insights into their robustness and providing useful guidelines for their tuning under practical scenarios.
Several open problems remain to be addressed. As previously mentioned, implementing the model-free Gauss-Newton method presents an interesting opportunity for future research. While this work focused primarily on direct data-driven policy gradient approaches, it would also be valuable to explore and compare indirect data-driven policy gradient methods, which integrate online parameter estimation with model-based gradient descent. A thorough comparison of direct and indirect PG methods in the presence of stochastic noise is an important topic for future study.
Acknowledgments
We would like to thank Nicolas Chatzikiriakos for his constructive discussion and valuable feedback. Bowen Song acknowledges the support of the International Max Planck Research School for Intelligent Systems (IMPRS-IS). Andrea Iannelli acknowledges the German Research Foundation (DFG) for support of this work under Germany’s Excellence Strategy - EXC 2075 – 390740016.
Appendix A Proof of Results in Section 2
A.1 Proof of Lemma 1
Proof 1
To prove Lemma 1, we introduce the kernel defined as $A(K^{\prime},K):=P^{K^{\prime}}_{K}-P_{K}$, where $P^{K^{\prime}}_{K}:=Q_{K^{\prime}}+(A+BK^{\prime})^{\top}P_{K}(A+BK^{\prime})$ with $K^{\prime},K\in\mathcal{S}$, which works as a bridge to connect $P_{K}$ and $P_{K^{\prime}}$. Then we have:
$\begin{split}A(K^{\prime},K)&=Q+K^{\prime\top}RK^{\prime}+(A+BK^{\prime})^{\top}P_{K}(A+BK^{\prime})-P_{K}\\ &=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &~~~~+(K^{\prime}-K)^{\top}\left((R+B^{\top}P_{K}B)K+B^{\top}P_{K}A\right)\\ &~~~~+\left((R+B^{\top}P_{K}B)K+B^{\top}P_{K}A\right)^{\top}(K^{\prime}-K)\\ &=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &=+(K^{\prime}-K)^{\top}E_{K}+E_{K}^{\top}(K^{\prime}-K).\end{split}$ (46)
A lower bound of (46) can be derived as:
$\begin{split}A(K^{\prime},K)&=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &~~~~+(K^{\prime}-K)^{\top}E_{K}+E_{K}^{\top}(K^{\prime}-K)\\ &=(\underbrace{K^{\prime}-K+(R+B^{\top}P_{K}B)^{-1}E_{K}}_{=:\mathcal{A}})^{\top}(R+B^{\top}P_{K}B)\mathcal{A}\\ &~~~~-E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K}\\ &\succeq-E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K},\end{split}$ (47)
with equality when $K^{\prime}=K-(R+B^{\top}P_{K}B)^{-1}E_{K}$. Then for $P_{K}$ and $P_{K^{\prime}}$, we have:
$P_{K}-P_{K^{\prime}}=\sum_{t=0}^{\infty}(A_{K}^{t})^{\top}A(K,K^{\prime})A_{K}^{t}.$ (48)
From (8) and (48), we set $K^{\prime}=K^{*}$ and then have:
[…]
From (11) and $\Sigma_{K}\succeq\Sigma_{w}\succ 0$, we can get the expression of $E_{K}$ as:
$E_{K}=\frac{1}{2}\nabla C(K)\Sigma_{K}^{-1}$
Then (49) can be further simplified as:
$\begin{split}C(K)-C(K^{*})&=\mathrm{Tr}\left(E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K}\Sigma_{K^{*}}\right)\\ &=\frac{1}{4}\mathrm{Tr}\left(\left(\Sigma_{K}\right)^{-1~\top}\nabla C(K)^{\top}(R+B^{\top}P_{K}B)^{-1}\nabla C(K)\left(\Sigma_{K}\right)^{-1}\Sigma_{K^{*}}\right)\\ &\leq\frac{1}{4}\lVert\Sigma_{K^{*}}\rVert\mathrm{Tr}\left(\nabla C(K)^{\top}(R+B^{\top}P_{K}B)^{-1}\nabla C(K)\right)\\ &=\frac{1}{4}\lVert\Sigma_{K^{*}}\rVert\lVert\Sigma_{w}^{-2}\rVert\lVert R^{-1}\rVert\lVert\nabla C(K)\rVert_{F}^{2}=\mu\lVert\nabla C(K)\rVert_{F}^{2}\end{split}$
This concludes the proof.
A.2 Proof of Lemma 2
Proof 2
From (48),
$\begin{split}&~~~~~C(K^{\prime})-C(K)\\ &=\mathrm{Tr}\left((P_{K^{\prime}}-P_{K})\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(\left(\sum_{t=0}^{\infty}(A_{K^{\prime}}^{t})^{\top}A({K^{\prime}},K)A_{K^{\prime}}^{t}\right)\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(\left[({K^{\prime}}-{K})^{\top}(R+B^{\top}P_{K}B)({K^{\prime}}-{K})+({K^{\prime}}-{K})^{\top}E_{K}+E_{K}^{\top}({K^{\prime}}-{K})\right]\Sigma_{K^{\prime}}\right)\end{split}$ (50)
The third equality is obtained by using (46). Then from (50), we can get:
$\left\lvert C(K^{\prime})-C(K)-2\mathrm{Tr}\left(({K^{\prime}}-{K})^{\top}E_{K}\Sigma_{K^{\prime}}\right)\right\rvert\leq\lVert\Sigma_{K^{\prime}}\rVert\left\lVert R+B^{\top}P_{K}B\right\rVert\left\lVert K^{\prime}-{K}\right\rVert_{F}$
This concludes the proof.
A.3 Proof of Lemma 3
Proof 3
From (8), we can derive the upper bound on the following matrices:
$\lVert P_{K}\rVert\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})},~~~\lVert\Sigma_{K}\rVert\leq\frac{C(K)}{\lambda_{1}(Q)}.$ (51)
Define a linear operator $\mathcal{T}_{K}$ as:
$\mathcal{T}_{K}(X):=\sum_{t=0}^{\infty}(A+BK)^{t}(X)(A+BK)^{t~\top}.$
Define the induced norm of $\mathcal{T}_{K}$ as:
$\lVert\mathcal{T}_{K}\rVert:=\mathop{sup}_{X}\frac{\lVert\mathcal{T}_{K}(X)\rVert}{\lVert X\rVert}.$
From [9, Lemma 17], we can derive the upper bound of $\lVert\mathcal{T}_{K}\rVert$ as
$\lVert\mathcal{T}_{K}\rVert\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}$
Define another linear operator $\mathcal{F}_{K}$ as:
$\mathcal{F}_{K}(X):=(A+BK)X(A+BK)^{\top}.$
Define the identity operator $\mathcal{I}(X):=X$ Then we have the following results from [9, Lemma 18]:
$\mathcal{T}_{K}=(\mathcal{I}-\mathcal{F}_{K})^{-1}$
By the definition of $\Sigma_{K}$
$\Sigma_{K}=\mathcal{T}_{K}(\Sigma_{w})=(\mathcal{I}-\mathcal{F}_{K})^{-1}(\Sigma_{w})$
From [9, Lemma 19]:
$\lVert\mathcal{F}_{K}-\mathcal{F}_{K^{\prime}}\rVert\leq 2\lVert A+BK\rVert\lVert B\rVert\lVert K-K^{\prime}\rVert+\lVert B\rVert^{2}\lVert K-K^{\prime}\rVert^{2}$
From [9, Lemma 20], if
$\lVert\mathcal{T}_{K}\rVert\lVert\mathcal{F}_{K}-\mathcal{F}_{K^{\prime}}\rVert\leq\frac{1}{2},$
then:
$\lVert(\mathcal{T}_{K}-\mathcal{T}_{K^{\prime}})(\Sigma_{w})\rVert\leq 2\lVert\mathcal{T}_{K}\rVert^{2}\lVert\mathcal{F}_{K}-\mathcal{F}_{K^{\prime}}\rVert\lVert\Sigma_{w}\rVert$
Then from [9, Lemma 16], we can prove the result of perturbation of $\Sigma_{K}$, as stated in Lemma 3
A.4 Proof of Lemma 4
Proof 4
When $\lVert K-K^{\prime}\rVert=\min\{h(C(K)),\lVert K\rVert\}$, by [9, Lemma 24], an upper bound on $\lVert P_{K^{\prime}}-P_{K}\rVert$ can be derived as:
$\begin{split}\lVert P_{K^{\prime}}-P_{K}\rVert\leq{6\left(\frac{C(K)}{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}\right)^{2}\left(2\lVert K\rVert^{2}\lVert R\rVert\lVert B\rVert+\lVert K\rVert\lVert R\rVert\right)}\lVert K-K^{\prime}\rVert\end{split}$ (52)
We aim to derive an upper bound on $\lVert K\rVert$ as a function of $C(K)$. First, we find the upper bound of the $C(K)-C(K^{*})$, we choose $K^{\prime}=K-(R+B^{\top}P_{K}B)^{-1}E_{K}$, from (47), we can get:
[…]
From (53), we can also derive the upper bound of $\lVert K\rVert$:
[…]
From (8) and (52), we can obtain
$\begin{split}&~~~|C(K^{\prime})-C(K)|=\mathrm{Tr}\left((P_{K^{\prime}}-P_{K})\Sigma_{w}\right)\leq\lVert P_{K^{\prime}}-P_{K}\rVert\mathrm{Tr}(\Sigma_{w})\\ &\leq\underbrace{6\left(\frac{C(K)}{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}\right)^{2}\left(2b^{2}_{K}(C(K))\lVert R\rVert\lVert B\rVert+b_{K}(C(K))\lVert R\rVert\right)\mathrm{Tr}(\Sigma_{w})}_{=:h_{C}(C(K))}\lVert K-K^{\prime}\rVert.\end{split}$ (55)
A.5 Proof of Lemma 5
Proof 5
From (11), we know that: $\nabla C(K)=2E_{K}\Sigma_{K}$, therefore:
$\begin{split}|\nabla C(K^{\prime})-\nabla C(K)|=2E_{K^{\prime}}\Sigma_{K^{\prime}}-2E_{K}\Sigma_{K}=2(E_{K^{\prime}}-E_{K})\Sigma_{K^{\prime}}+2E_{K}(\Sigma_{K^{\prime}}-\Sigma_{K})\end{split}$
For the first term, we aim to investigate the upper bound of $E_{K}$. From (53), we know that:
$\begin{split}\mathrm{Tr}\left(E_{K}^{\top}E_{K}\right)\leq\frac{(C(K)-C(K^{*}))\lVert R+B^{\top}P_{K}B\rVert}{\sigma_{1}(\Sigma_{w})}\\ \leq\frac{(C(K)-C(K^{*}))}{\sigma_{1}(\Sigma_{w})}\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{w})}}\right).\end{split}$
Then, an upper-bound on the gradient is given as:
$\begin{split}\lVert\nabla C(K)\rVert&\leq\sqrt{4\lVert\Sigma_{K}\rVert^{2}\mathrm{Tr}\left(E_{K}^{\top}E_{K}\right)}\\ &\leq\underbrace{\sqrt{4\left(\frac{C(K)}{\lambda_{1}(Q)}\right)^{2}\frac{(C(K)-C(K^{*}))}{\sigma_{1}(\Sigma_{w})}\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{w})}}\right)}}_{=:b_{\nabla}(C(K))}.\end{split}$ (56)
[…]
$\begin{split}&~~~~\lVert 2E_{K}(\Sigma_{K}-\Sigma_{K^{\prime}})\rVert\\ &\leq\underbrace{2\sqrt{\frac{(C(K)-C(K^{*}))}{\sigma_{1}(\Sigma_{w})}\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{w})}}\right)}h_{\Sigma}{(C(K))}}_{=:\alpha_{1}(C(K))}\lVert K-K^{\prime}\rVert.\end{split}$ (57)
For the term $E_{K^{\prime}}-E_{K}$, we need to find the upper bound of $P_{K^{\prime}}-P_{K}$: by (52), for $\lVert K-K^{\prime}\rVert=\min\{h\left(C(K),K\right),\lVert K\rVert\}$, we have
$\begin{split}\lVert P_{K^{\prime}}-P_{K}\rVert\leq\underbrace{6\left(\frac{C(K)}{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}\right)^{2}\left(2b^{2}_{K}(C(K))\lVert R\rVert\lVert B\rVert+b_{K}(C(K))\lVert R\rVert\right)}_{=:\alpha_{2}(C(K))}\lVert K-K^{\prime}\rVert.\end{split}$
From Lemma 3:
$\lVert\Sigma_{K}-\Sigma_{K^{\prime}}\rVert\leq\frac{C(K)}{\lambda_{1}(Q)}.$
Then we know:
$\lVert\Sigma_{K^{\prime}}\rVert\leq\lVert\Sigma_{K}\rVert+\lVert\Sigma_{K}-\Sigma_{K^{\prime}}\rVert\leq\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}.$
Therefore
$\begin{split}E_{K^{\prime}}-E_{K}&=R(K^{\prime}-K)+B^{\top}(P_{K^{\prime}}-P_{K})A\\ &~~+B^{\top}(P_{K^{\prime}}-P_{K})BK^{\prime}+B^{\top}P_{K}B(K^{\prime}-K).\end{split}$
With assumption $\lVert K^{\prime}-K\rVert\leq\lVert K\rVert$, we can get $\lVert K^{\prime}\rVert\leq 2\lVert K\rVert$. Combining all the equations, an upper bound on $2(E_{K^{\prime}}-E_{K})\Sigma_{K^{\prime}}$ can be derived as:
$\begin{split}&\lVert 2(E_{K^{\prime}}-E_{K})\Sigma_{K^{\prime}}\rVert\leq 2\lVert E_{K^{\prime}}-E_{K}\rVert\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\\ &\leq 2\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{0})}}\right)\lVert K-K^{\prime}\rVert\\ &~~~+2\alpha_{2}(C(K))\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\left(\lVert B\rVert\lVert A\rVert+\lVert B\rVert^{2}\lVert K\rVert\right)\lVert K-K^{\prime}\rVert\\ &\leq\alpha_{3}(C(K))\lVert K-K^{\prime}\rVert\end{split}$ (58)
with
$\alpha_{3}(C(K)):=\left[\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{0})}}+\alpha_{2}(C(K))\left(\lVert B\rVert\lVert A\rVert+b_{K}(C(K))\lVert B\rVert^{2}\right)\right].$
Combing (57) and (58), we have:
$\lVert\nabla C(K^{\prime})-\nabla C(K)\rVert\leq\underbrace{(\alpha_{1}(C(K))+\alpha_{3}(C(K)))}_{=:h_{\nabla}(C(K))}\lVert K-K^{\prime}\rVert$ (59)
when $\lVert K-K^{\prime}\rVert=\min\{h(C(K)),\lVert K\rVert\}$.
Appendix B Proof of Results in Section 3
B.1 Proof of Theorem 1
Proof 6
Using Lemma 2 and following the step by [9, Lemma 21], if
$\begin{split}\eta_{i}\leq\frac{1}{32}\min\left\{\left(\frac{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}{C(K_{i})}\right)^{2}\frac{1}{\lVert B\rVert\lVert\nabla C(K_{i})\rVert(1+\lVert A+BK_{i}\rVert)},\right.\\ \left.\frac{\lambda_{1}(Q)}{2C(K_{i})\lVert R+B^{\top}P_{K_{i}}B\rVert}\right\},\end{split}$ (60)
then,
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K}^{*}\rVert}\right)(C(K_{i})-C(K^{*})).$ (61)
Based on the upper bounds on $\lVert\nabla C(K)\rVert$, we can derive the following from (60):
$\eta_{i}\leq\underbrace{\frac{1}{32}\min\left\{\left(\frac{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}{C(K_{i})}\right)^{2}\frac{1}{2\lVert B\rVert b_{\nabla}(C(K_{i}))},\frac{\lambda_{1}(Q)}{2C(K_{i})\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}{(\Sigma_{w})}}\right)}\right\}}_{=:h_{\mathrm{PGD}}(C(K))},$ (62)
where $b_{\nabla}$ is defined in (56). From (61), we know that $C(K_{i+1})\leq C(K_{i})$. Then we have $h_{\mathrm{PGD}}(C(K_{i+1}))\geq h_{\mathrm{PGD}}(C(K_{i}))$. The remainder of the proof proceeds by setting $\eta_{i}=\eta_{0},\forall i\in\mathbb{Z}_{+}$, which represents the lower bound of the convergence rate, and then following the steps outlined in [9, Lemma 22].
B.2 Update of Natural Policy Gradient
Proof 7
Vectorizing the gain $K$ as $\bar{K}$:
$K\in\mathbb{R}^{n_{x}\times n_{u}}=\left[\begin{array}[]{cc}k_{1}\\ k_{2}\\ ...\\ k_{n_{x}}\end{array}\right]\rightarrow[k_{1},k_{2},...,k_{n_{x}}]^{\top}=:\bar{K}\in\mathbb{R}^{n_{x}n_{u}}$
Then the input $u_{t}=Kx_{t}$ can be expressed as:
$u_{t}=\underbrace{\left[\begin{array}[]{cccc}x_{t}^{\top}&0&...&0\\ 0&x_{t}^{\top}&...&...\\ ...&...&...&0\\ 0,&...&0&x_{t}^{\top}\end{array}\right]}_{=:X_{t}\in\mathbb{R}^{n_{u}\times n_{x}n_{u}}}\left[\begin{array}[]{cc}k_{1}^{\top}\\ k_{2}^{\top}\\ ...\\ k_{n_{x}}^{\top}\end{array}\right]=X_{t}\bar{K}=Kx_{t}$ (63)
We parameterize the linear policy with additive Gaussian noise with
$u_{t}=\pi_{\bar{K}}(x_{t},u_{t})=\mathcal{N}(X_{t}K,\alpha^{2}I)$
Based on the PDF of Normal distribution:
$\pi_{\bar{K}}(u_{t}|x_{t})=\frac{1}{\sqrt{(2\pi)^{n_{x}n_{u}}\alpha^{2}}}\mathop{exp}\left[-\frac{1}{2\alpha^{2}}(u_{t}-X_{t}\bar{K})^{\top}(u_{t}-X_{t}\bar{K})\right]$
The derivative with respect to $\bar{K}$:
$\begin{split}\mathbb{E}\left[\nabla\mathop{log}\pi_{\bar{K}}(u_{t}|x_{t})\nabla\mathop{log}\pi_{\bar{K}}(u_{t}|x_{t})^{\top}\right]&=\mathbb{E}\left[\frac{X_{t}^{\top}(u_{t}-X_{t}\bar{K})(u_{t}-X_{t}\bar{K})^{\top}X_{t}}{\alpha^{4}}\right]\\ &=\left[\begin{array}[]{cccc}\Sigma_{K}&0&...&0\\ 0&\Sigma_{K}&...&...\\ ...&...&...&0\\ 0,&...&0&\Sigma_{K}\end{array}\right]\in\mathbb{R}^{n_{x}n_{u}\times n_{x}n_{u}}\end{split}$
Then $G_{\bar{K}}^{-1}$ are given as:
$G_{\bar{K}}^{-1}=\left[\begin{array}[]{cccc}\Sigma_{K}^{-1}&0&...&0\\ 0&\Sigma_{K}^{-1}&...&...\\ ...&...&...&0\\ 0,&...&0&\Sigma_{K}^{-1}\end{array}\right].$
With the inverse matrix transformation as (63), we can get:
$K_{t+1}=K_{t}-\eta\nabla C(K)\Sigma_{K}^{-1}$
Appendix C Proof of Results in Section 4
C.1 Approximating $C(K),\Sigma_{K}$ with finite horizon
We aim to show that, for $K\in\mathcal{S}$, it is possible to approximate both $C(K)$ and $\Sigma_{K}$ with any desired accuracy:
Lemma 6 (Approximating $C(K)$ and $\Sigma_{K}$ with finite horizon)
For any $K\in\mathcal{S}$, let
$\Sigma_{K}^{(l)}:=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{x_{0},w_{t}}\left[x_{t}x_{t}^{\top}\right]~\mathrm{and}~C^{(l)}(K):=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{x_{0},w_{t}}\left[x_{t}^{\top}Q_{K}x_{t}\right].$ (64)
Given any desired accuracy $\epsilon_{\Sigma},\epsilon_{C}>0$, if
$l\geq\frac{2C(K)}{\epsilon_{\Sigma}}{\left(\frac{\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{1}{\lambda_{1}(Q)}\right)},$
then
$\lVert\Sigma_{K}^{(l)}-\Sigma_{K}\rVert\leq\epsilon_{\Sigma}.$
If
$l\geq\frac{2C(K)}{\epsilon_{C}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right),$ (65)
then
$\lVert C^{(l)}(K)-C(K)\rVert\leq\epsilon_{C}.$
Proof 8
We define $\tilde{P}_{K}^{w}$ as the solution to the following equation:
$\tilde{P}_{K}^{w}=A_{K}P_{K}^{w}A_{K}^{\top}+A_{K}\tilde{P}_{K}^{w}A_{K}^{\top}$
and the operator $\mathcal{C}_{Y}(X):=\sum_{t=0}^{+\infty}Y^{t}XY^{t~\top}$ with $X\succ 0$ and $\lVert Y\rVert\leq 1$ . From the definition of $\Sigma_{K}^{(l)}$:
[…]
We take the norm for both sides:
$\begin{split}\lVert\Sigma_{K}^{(l)}-\Sigma_{K}\rVert&\leq\frac{2\lVert\mathcal{C}_{A_{K}}(\Sigma_{0}-P_{K})\rVert+2\lVert\Sigma_{K}\rVert}{l}\\ &\leq\frac{2}{l}{\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right)}=:\epsilon_{\Sigma}.\end{split}$
Then for any desired accuracy $\lVert\Sigma_{K}^{(l)}-\Sigma_{K}\rVert\leq\epsilon_{\Sigma}$, we can choose a rollout length
$l\geq\frac{2C(K)}{\epsilon_{\Sigma}}{\left(\frac{\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{1}{\lambda_{1}(Q)}\right)}$
to satisfy the accuracy.
[…]
We take the absolute value for both sides:
[…]
Then for any desired accuracy $\lVert C^{(l)}(K)-C(K)\rVert\leq\epsilon_{C}$, we can choose a rollout length:
$l\geq\frac{2C(K)}{\epsilon_{C}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right).$ (66)
C.2 Smoothing and the gradient descent analysis
The algorithm performs gradient descent on the following smoothing function:
$C_{r}(K):=\mathbb{E}_{U\sim\mathbb{B}_{r}}\left[C(K+U)U\right],$ (67)
where $\mathbb{B}_{r}$ is the uniform distribution over all matrices with Frobenius norm less than $r$ (the entire ball)
Lemma 7 (Gradient of smoothing function)
$\nabla C_{r}(K)=\frac{n_{x}n_{u}}{r^{2}}\mathbb{E}_{U\sim\mathbb{S}_{r}}\left[C(K+U)U\right],$ (68)
where $\mathbb{S}_{r}$ is the uniform distribution over all matrices with Frobenius norm $r$ (the sphere)
The proof of Lemma 7 is given in [9, Supplementary Material, Lemma 26].
C.3 Estimating $\nabla C(K)$ with finitely many infinite-horizon Rollouts
In this section, we aim to bound the difference between $\nabla C(K)-\bar{\nabla}C(K)$, where
$\bar{\nabla}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}C(K+U_{k})U_{k},$ (69)
and $U_{i}$ is uniformly distributed and $\lVert U_{i}\rVert_{F}=r$.
Lemma 8
Given an accuracy $\epsilon_{nr}$ and probability $\delta_{n}$, which can be expressed by $\epsilon_{r}$ and $\epsilon_{n}$ ($\epsilon_{nr}=\epsilon_{r}+\epsilon_{n}$), there exist an radius $r_{\max}$ which is a function of $\epsilon_{r}$ and the number of rollouts $N_{1}$ which is a function of $\epsilon_{r},\epsilon_{n}$ such that when $r\leq r_{\max}$ and $n\geq N_{1}$,
$\mathbb{P}\left\{\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert\leq\epsilon_{nr}\right\}\geq 1-\delta_{n}.$ (70)
Proof 9
First, we bridge $\bar{\nabla}C(K)$ and $\nabla C(K)$ with $\nabla C_{r}(K)$,
$\bar{\nabla}C(K)-\nabla C(K)=(\nabla C_{r}(K)-\nabla C(K))+(\bar{\nabla}C(K)-\nabla C_{r}(K)).$
For the first term, let $r_{\max}\leq\min\{h(C(K)),\lVert K\rVert\}$,we have $C(K+U_{k})-C(K)\leq h_{C}(C(K))\lVert U_{k}\rVert_{F}$, Then we can get $|C(K+U_{k})|\leq C(K)+h_{C}(C(K))\lVert U_{k}\rVert_{F}$. From Lemma 5, if
$r_{\max}\leq\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{r}}{h_{\nabla}(C(K))}\},$ (71)
then
$\lVert\nabla C(K+U_{k})-{\nabla}C(K)\rVert\leq{\epsilon_{r}}.$
Since $\nabla C_{r}(K)$ is the expectation of $\nabla C(K+U_{k})$, i.e.
$\nabla C_{r}(K)=\mathbb{E}_{U\sim\mathbb{S}_{r}}\left[\nabla C(K+U)\right],$
by the triangle inequality, we have
$\lVert\nabla C_{r}(K)-\nabla C(K)\rVert\leq{\epsilon_{r}}.$ (72)
For the second term $\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert$, we define the sample $Z_{k}:=\frac{n_{x}n_{u}}{r^{2}}C(K+U_{k})U_{k}$
$\lVert Z_{k}\rVert\leq\frac{n_{x}n_{u}}{r^{2}}C(K+U_{k})\lVert U_{k}\rVert\leq\frac{n_{x}n_{u}C(K+U_{k})}{r}\leq\frac{n_{x}n_{u}\left(C(K)+rh_{C}(C(K))\right)}{r}$
We define $Z:=\nabla C_{r}(K)$, together with (72), an upper bounded can be derived as:
$\lVert Z\rVert=\lVert\nabla C_{r}(K)\rVert\leq\lVert\nabla C(K)\rVert+{\epsilon_{r}}\leq{\epsilon_{r}}+b_{\nabla}(C(K)).$
Then an upper bound on $\lVert Z_{k}-Z\rVert$ as:
$\lVert Z_{k}-Z\rVert\leq\lVert Z_{k}\rVert+\lVert Z\rVert\leq\underbrace{\frac{n_{x}n_{u}\left(C(K)+rh_{C}(C(K))\right)}{r}+\epsilon_{r}+b_{\nabla}(C(K))}_{=:\alpha_{4}(C(K),\epsilon_{r})}$ (73)
Further, we have:
$\begin{split}&~~~~\lVert\mathbb{E}(Z_{k}Z_{k}^{\top})-ZZ^{\top}\rVert=\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})-Z^{\top}Z\rVert\\ &\leq\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})\rVert_{F}+\lVert Z^{\top}Z\rVert_{F}\leq\max_{Z_{k}}(\lVert Z_{k}\rVert_{F})^{2}+\lVert Z\rVert^{2}_{F}\\ &\leq\underbrace{\max\{n_{x},n_{u}\}^{2}\left(\frac{n_{x}n_{u}\left(C(K)+rh_{C}(C(K))\right)}{r}\right)^{2}+\left(\epsilon_{r}+b_{\nabla}(C(K))\right)^{2}}_{=:\alpha_{5}(C(K),\epsilon_{r})}\end{split}$ (74)
We aim to bound $\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert\leq\epsilon_{n}$, so the number of rollouts must be greater or equal to $N_{1}$ (from Matrix Bernstein Inequality, [10, Lemma B.5])
$N_{1}:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{n}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{r})+\frac{\alpha_{5}(C(K),\epsilon_{r})\epsilon_{n}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta}\right],$ (75)
where $\alpha_{4}$ and $\alpha_{5}$ are defined in (73) and (74), respectively. Then we have
$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert_{F}\leq\epsilon_{n}\right]\geq 1-\delta_{n}.$
We sum up all the terms and get the following: if $r\geq r_{\max}$ and ${n}\geq N$, where $r_{min}$ and $N_{1}$ are defined in (71) and (75), we have
$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert_{F}\leq\epsilon_{nr}\right]\geq 1-\delta_{n}.$ (76)
C.4 boundedness of state $\lVert x_{t}\rVert$
The noise $w_{t}$ is a Gaussian-distributed vector with zero mean and covariance $\Sigma_{w}$. Let $\Sigma_{w}=V\Lambda V^{\top}$ be the eigenvalue decomposition, where $\Lambda$ is a diagonal matrix with diagonal entries $\sigma_{1},...,\sigma_{n_{x}}$. Then we have the vector $W_{t}=V^{-1}w_{t}$, with zero mean and variance $\Lambda$. By the Chebyshev’s inequality [52, Corollary 1.2.5], we have
$\mathbb{P}\left\{\lVert w_{t}\rVert\geq\bar{w}\right\}=\mathbb{P}\left\{\lVert W_{t}\rVert\geq\bar{w}\right\}\leq\frac{\sum_{k=1}^{n_{x}}\sigma_{k}}{\bar{w}^{2}}=\frac{\mathrm{Tr}(\Sigma_{w})}{\bar{w}^{2}}=:\delta_{w}({\bar{w}})$ (77)
Based on (77), we can derive the following Lemma:
Lemma 9 (Bound of state $\lVert x_{t}\rVert$)
For any initial state $\lVert x_{0}\rVert\leq L_{0}$ and $K\in\mathcal{S}$, given a desired probability $\delta_{x}$, the state response $\lVert x_{t}\rVert$ expressed as (3) is bounded by $\bar{L}:=L_{0}+t\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/t}}$ with probability at least $1-\delta_{x}$.
Proof 10
From (3), we can get:
$\begin{split}\lVert x_{t}\rVert&\leq\lVert(A_{K})^{t}x_{0}^{(i)}\rVert+\left\lVert\sum_{k=1}^{t}(A_{K})^{t-k}w_{k-1}\right\rVert\\ &\leq\lVert x_{0}\rVert+\sum_{k=1}^{t}\lVert(A_{K})^{t-k}\rVert\left\lVert w_{k-1}\right\rVert\leq L_{0}+\sum_{k=0}^{t-1}\left\lVert w_{k}\right\rVert\end{split}$
If for all $t\in[0,t-1]$$\lVert w_{t}\rVert\leq\bar{w}$, then $\lVert x_{t}\rVert\leq L_{0}+t\bar{w}=\bar{L}$. Then the probability of $\lVert x_{t}\rVert\leq\bar{L}$ is at least $1-\delta=(1-\delta_{w}(\bar{w}))^{t}$. Together with (77), we can then derive $\bar{w}=\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/t}}$. We conclude the proof.
C.5 Proof of Theorem 4
We aim to bound the difference between the estimates from the Algorithm 1, expressed by
$\hat{\nabla}C(K)=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{\bar{K}_{k}}U_{k}$
and the true gradient $\nabla C(K)$.
Proof 11
We bridge this two terms with $\nabla^{\prime}C(K)$ and $\bar{\nabla}C(K))$, where $\bar{\nabla}C(K))$ is defined in Lemma 8 and $\nabla^{\prime}C(K)$ is defined as:
$\nabla^{\prime}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}C^{(l)}(K+U_{k})U_{k},$
where $C^{(l)}(K)$ is defined in Lemma 6. The relation between the estimates and the true gradient is given as:
$\hat{\nabla}C(K)-\nabla C(K)=(\hat{\nabla}C(K)-\nabla^{\prime}C(K))+(\nabla^{\prime}C(K)-\bar{\nabla}C(K))+(\bar{\nabla}C(K)-\nabla C(K))$ (78)
The third term is analyzed in Lemma 8. We choose $\epsilon_{nr}:=\epsilon_{n}+\epsilon_{r}$ and $\delta_{n}$, then we can find the exploration radius $r_{\max}$ and $N_{1}$ as:
$r_{\max}$ $(C(K),\epsilon_{r})=\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{r}}{h_{\nabla}(C(K))}\},$ (79a)
$N_{1}$ $=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{n}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{r})+\frac{\alpha_{5}(C(K),\epsilon_{r})\epsilon_{n}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{n}}\right]$ (79b)
Then we have
$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert_{F}\leq\epsilon_{nr}\right]\geq 1-\delta_{n}.$ (80)
For the second term, as stated in Lemma 6, for any desired accuracy $\epsilon_{l}$, we can choose $l$ such that:
$l\geq\frac{2n_{x}n_{u}(C(K)+rh_{C}C(K))^{2}}{\epsilon_{l}r\lambda_{1}(\Sigma_{w})}\left(\frac{\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{(C(K)+rh_{C}C(K))}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{1}{\lambda_{1}(Q)}\right),$ (81)
then,
$|C^{(l)}(\bar{K}_{k})-C(\bar{K}_{k})|\leq\frac{r\epsilon_{l}}{n_{x}n_{u}},\quad\forall k\in[1,n].$ (82)
By the triangle inequality and $\lVert U_{k}\rVert_{F}=r$, we can get:
$\lVert\nabla^{\prime}C(K)-\bar{\nabla}C(K)\rVert\leq\left\lVert\frac{1}{n}\sum_{k=1}^{n}\frac{n_{x}n_{u}}{r^{2}}\left[C^{(l)}(\bar{K}_{k})-C(\bar{K}_{k})\right]U_{k}\right\rVert=\epsilon_{l}$ (83)
For the first term, from Lemma 9, we know that for any selected accuracy $\delta_{x}$, we have
$\lVert x_{t}\rVert\leq\bar{L}=L_{0}+t\bar{w}=L_{0}+t\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/t}}$ (84)
with probability at least $1-\delta_{x}$. Given the probability $\delta_{x}$, the upper bound of the sample $\hat{C}_{\bar{K}_{k}}$ is given by:
$\begin{split}&~~~~\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~\top}(Q+\bar{K}^{\top}_{k}R\bar{K}_{k})x_{t}^{(k)}\\ &\leq\frac{1}{l}\sum_{t=0}^{l-1}\lVert Q_{\bar{K}_{k}}\rVert\lVert x_{t}^{(k)}\rVert^{2}\leq(\lVert Q_{\bar{K}_{k}}\rVert\max_{t\in[0,l-1]}\lVert x_{t}^{(k)}\rVert^{2})\leq\frac{C(\bar{K}_{k})}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\bar{L}\right)^{2}\\ &\leq\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/l}}\right)^{2}=:\bar{C}\end{split}$ (85)
Each sample $Z_{k}:=\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{\bar{K}_{k}}U_{k}$ has bounded norm as:
$\lVert Z_{k}\rVert\leq\frac{n_{x}n_{u}}{r}\bar{C}=:\alpha_{6}(C(K),\delta_{x}).$ (86)
Then we have for $Z:=\mathbb{E}[\hat{\nabla}C(K)]=\nabla C^{\prime}(K)$, with (80) and (83),
$\lVert Z\rVert\leq\epsilon_{l}+\lVert\bar{\nabla}C(K)\rVert\leq\epsilon_{l}+\epsilon_{nr}+\lVert\nabla{C}(K)\rVert\leq\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))$ (87)
Then we can derive the bound on $\lVert Z_{k}-Z\rVert$:
$\lVert Z_{k}-Z\rVert\leq\lVert Z\rVert+\lVert Z_{k}\rVert\leq\underbrace{\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))+\alpha_{6}(C(K),\delta_{x})}_{=:\alpha_{7}(C(K),\delta_{x})}$ (88)
Similarly, the bound on $\lVert\mathbb{E}(Z_{k}Z_{k}^{\top})-ZZ^{\top}\rVert$ can be derived as:
$\begin{split}\lVert\mathbb{E}(Z_{k}Z_{k}^{\top})-ZZ^{\top}\rVert&\leq\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})\rVert_{F}+\lVert Z^{\top}Z\rVert_{F}\leq\max_{Z_{k}}(\lVert Z_{k}\rVert_{F})^{2}+\lVert Z\rVert^{2}_{F}\\ &\leq\underbrace{\max\{n_{x},n_{u}\}^{2}\alpha^{2}_{6}(C(K),\epsilon_{x})+\left(\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))\right)^{2}}_{=:\alpha_{8}(C(K),\delta_{x})}\end{split}$ (89)
Given an accuracy $\epsilon_{d}$ and $\delta_{d}$, we can use the matrix concentration inequality [10, Lemma B2.5],
$N_{2}=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{8}(C(K),\delta_{x})\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right]$ (90)
Then we have
$\mathbb{P}\left[\lVert\hat{\nabla}C(K)-\nabla^{\prime}C(K)\rVert_{F}\leq\epsilon_{d}\right]\geq 1-\delta_{d}.$ (91)
Now, we can summarize all the terms. Given an arbitrary tolerance $\epsilon$, which can be expressed by $\epsilon_{d},\epsilon_{l},\epsilon_{n},\epsilon_{r}$, with $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$, and an arbitrary probability $\delta$, which can be expressed by $\delta_{x},\delta_{d},\delta_{n}$ with $1-\delta=(1-\delta_{d})(1-\delta_{n})(1-\delta_{x})$, suppose the exploration radius is chosen as $r\leq r_{\max}$ with
$r_{\max}=\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{r}}{h_{\nabla}(C(K))}\},$ (92)
where $h_{\nabla}$ is defined in (59), the rollout length is chosen as: $l\geq l_{\min}$
$l_{\min}=\frac{2n_{x}n_{u}(C(K)+rh_{C}C(K))^{2}}{\epsilon_{l}r\lambda_{1}(\Sigma_{w})}\left(\frac{\lVert\Sigma_{0}\rVert\lambda_{1}(\Sigma_{w})+C(K)+rh_{C}C(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{1}{\lambda_{1}(Q)}\right),$ (93)
and the number of rollouts is chosen as: $n\geq\max\{N_{1},N_{2}\}$ with
$N_{1}$ $:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{n}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{r})+\frac{\alpha_{5}(C(K),\epsilon_{r})\epsilon_{n}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{n}}\right],$ (94a)
$N_{2}$ $:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{8}(C(K),\delta_{x})\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right],$ (94b)
where $\alpha_{4},\alpha_{5},\alpha_{7}$ and $\alpha_{8}$ are defined in (73), (74), (88) and (89) respectively. Then, we have
$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq 1-\delta.$
C.6 Proof of Theorem 5
Proof 12
First, we also bridge the estimates $\hat{\Sigma}_{K}$ and $\Sigma_{K}$ with the following equation:
$\hat{\Sigma}_{K}-\Sigma_{K}=(\hat{\Sigma}_{K}-\tilde{\Sigma}^{(l)}_{K})+(\tilde{\Sigma}^{(l)}_{K}-\tilde{\Sigma}_{K})+(\tilde{\Sigma}_{K}-\Sigma_{K}),$ (95)
with $\tilde{\Sigma}^{(l)}_{K}:=\frac{1}{n}\sum_{i=1}^{n}\Sigma^{(l)}_{K+U_{i}}$ and $\tilde{\Sigma}_{K}:=\frac{1}{n}\sum_{i=1}^{n}\Sigma_{K+U_{i}}$.
For the last term, we aim to bound $\lVert\tilde{\Sigma}_{K}-\Sigma_{K}\rVert$. From Lemma 3, if we choose $r\leq r^{\prime}_{\max}$ as $r^{\prime}_{\max}:=\frac{\epsilon^{\prime}_{r}}{b_{\nabla}(C(K))}$, where $b_{\nabla}$ is defined in (56), then $\lVert\tilde{\Sigma}_{K}-\Sigma_{K}\rVert\leq{\epsilon^{\prime}_{r}}$.
For the second term, from Lemma 6, if we choose $l\geq l^{\prime}_{\min}$ with
$l^{\prime}_{\min}:=\frac{2C(K)}{\epsilon_{C}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right),$
then $\lVert\tilde{\Sigma}^{(l)}_{K}-\tilde{\Sigma}_{K}\rVert\leq\epsilon^{\prime}_{l}$.
Let us turn to the first term, from Lemma 9,
$\lVert\underbrace{\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(i)}x_{t}^{(i)~\top}}_{=:\tilde{Z}_{k}}\rVert\leq\max_{i\in[0,l-1]}\lVert x_{t}^{(i)}\rVert^{2}$
For any desired probability $\delta^{\prime}_{x}$, $\lVert x_{t}^{(k)}\rVert$ is upper-bounded by $\bar{L}^{\prime}$ which is defined as $\bar{L}^{\prime}(\delta^{\prime}_{x}):=\left(L_{0}+t\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta^{\prime}_{x})^{1/t}}\right)$(see Lemma 9). Because of the definition of $\tilde{\Sigma}^{(l)}_{K}$, we have:
$\lVert\tilde{\Sigma}^{(l)}_{K}\rVert\leq\frac{1}{n}\sum_{k=1}^{n}\frac{C(K+U_{k})}{\lambda_{1}(Q)}$
For $r\leq\min\{h(C(K)),\lVert K\rVert\}$, we have $C(K+U_{i})\leq C(K)+rh_{C}(C(K))$, then we can get:
$\lVert\tilde{\Sigma}^{(l)}_{K}\rVert\leq\frac{1}{n}\sum_{k=1}^{n}\frac{C(K+U_{k})}{\lambda_{1}(Q)}\leq\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(Q)}$
Now an upper bound on $\lVert\tilde{Z}_{k}-\tilde{Z}\rVert$ is given as:
$\lVert\tilde{Z}_{k}-\tilde{Z}\rVert\leq\lVert\tilde{Z}_{k}\rVert+\lVert\tilde{Z}\rVert\leq\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(Q)}+(\bar{L}^{\prime}(\delta^{\prime}_{x}))^{2}=:\alpha_{9}(C(K),\delta^{\prime}_{x})$ (96)
Then an upper bound on $\lVert\mathbb{E}(\tilde{Z}_{k}\tilde{Z}_{k}^{\top})-\tilde{Z}\tilde{Z}^{\top}\rVert$ is given as:
$\begin{split}\lVert\mathbb{E}(\tilde{Z}_{k}\tilde{Z}_{k}^{\top})-\tilde{Z}\tilde{Z}^{\top}\rVert&\leq\lVert\mathbb{E}(\tilde{Z}_{k}^{\top}\tilde{Z}_{k})\rVert_{F}+\lVert\tilde{Z}^{\top}\tilde{Z}\rVert_{F}\\ &\leq\max_{\tilde{Z}_{k}}(\lVert\tilde{Z}_{k}\rVert_{F})^{2}+\lVert\tilde{Z}\rVert^{2}_{F}\\ &\leq\underbrace{n_{x}^{2}\left[(\bar{L}^{\prime}(\delta^{\prime}_{x}))^{2}+\left(\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(Q)}\right)^{2}\right]}_{=:\alpha_{10}(C(K),\delta^{\prime}_{x})}\end{split}$ (97)
With Matrix Bernstein Inequality [10, Lemma B2.5], given a small tolerance $\epsilon^{\prime}_{n}$} and small probability $\delta^{\prime}_{n}$, if the number of rollouts $n\geq n_{\min}^{\prime}$ with
$n_{\min}^{\prime}:=\frac{2n_{x}}{(\epsilon^{\prime}_{n})^{2}}\left(\alpha_{10}(C(K),\delta^{\prime}_{x})+\frac{\alpha_{9}(C(K),\delta^{\prime}_{x})\epsilon^{\prime}_{n}}{3\sqrt{n_{x}}}\right)\log\left[\frac{2n_{x}}{\delta^{\prime}_{n}}\right]$
where $\alpha_{9}$ and $\alpha_{10}$ are defined in (96) and (97). Then we have
$\mathbb{P}\left[\lVert\hat{\Sigma}_{K}-\tilde{\Sigma}^{(l)}_{K}\rVert\leq\epsilon^{\prime}_{n}\right]\geq 1-\delta^{\prime}_{n}$
We sum all terms, we can derive that given an arbitrary tolerance $\epsilon^{\prime}$, which can be expressed by $\epsilon^{\prime}_{n},\epsilon^{\prime}_{l},\epsilon^{\prime}_{r}$, with $\epsilon=\epsilon^{\prime}_{n}+\epsilon^{\prime}_{l}+\epsilon^{\prime}_{r}$, and an arbitrary probability $\delta^{\prime}$, which can be expressed by $\delta^{\prime}_{x},\delta^{\prime}_{n}$ with $1-\delta^{\prime}=(1-\delta_{n}^{\prime})(1-\delta_{x}^{\prime})$, suppose the exploration radius is chosen as $r\leq r^{\prime}_{\max}$ with
$r^{\prime}_{\max}:=\min\left\{h(C(K)),\lVert K\rVert,\frac{\epsilon^{\prime}_{r}}{b_{\nabla}(C(K))}\right\}$ (98)
the rollout length is chosen as $l\geq l^{\prime}_{\min}$ with
$l^{\prime}_{\min}:=\frac{2C(K)}{\epsilon^{\prime}_{l}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right)$ (99)
and the number of rollouts is chosen as $n\geq n_{\min}^{\prime}$ with
$n_{\min}^{\prime}:=\frac{2n_{x}}{(\epsilon^{\prime}_{n})^{2}}\left(\alpha_{10}(C(K),\delta^{\prime}_{x})+\frac{\alpha_{9}(C(K),\delta^{\prime}_{x})\epsilon^{\prime}_{n}}{3\sqrt{n_{x}}}\right)\log\left[\frac{2n_{x}}{\delta^{\prime}_{n}}\right]$ (100)
then we have,
$\mathbb{P}\left\{\lVert\hat{\Sigma}_{K}-\Sigma_{K}\rVert\leq\epsilon^{\prime}\right\}\geq 1-\delta^{\prime}.$
C.7 Proof of Theorem 6
Proof 13
In this subsection, we aim to prove that the introduction of the baseline function reduces the sample complexity. From the proof of Theorem 4, the required sample $n_{\min}$ is determined by the upper bound of the sample $\lVert Z_{k}\rVert$. Using (85), we know that this upper bound is probabilistically bounded as follows:
$\begin{split}\lVert\hat{C}_{\bar{K}_{k}}-\hat{b}_{s}(x_{0})\rVert&=\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})+b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\\ &\leq\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})\rVert+\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\end{split}$ (101)
Since $\hat{b}_{s}(x_{0})>0$, the term $\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})\rVert$ can be bounded as follows. If ${b}_{s}(x_{0})\leq\hat{C}_{\bar{K}_{k}}$, then bounded by $\bar{C}-{b}_{s}(x_{0})$ where $\bar{C}\geq\hat{C}_{\bar{K}_{k}}$ for all $k$ and $\bar{C}$ was introduced in (85). Otherwise, if ${b}_{s}(x_{0})\leq\hat{C}_{\bar{K}_{k}}$, and by ${b}_{s}(x_{0})$, if ${b}_{s}(x_{0})>\hat{C}_{\bar{K}_{k}}$. Therefore, we have:
$\begin{split}\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})\rVert\leq\max\{\hat{C}_{\bar{K}_{k}}-{b}_{s}(x_{0}),{b}_{s}(x_{0})\}\leq\max\{\bar{C}-{b}_{s}(x_{0}),{b}_{s}(x_{0})\}\end{split}$ (102)
From the definition of $b_{s}(x_{0})$, we know:
$b_{s}(x_{0})\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}(\lVert x_{0}\rVert+(l-1)\lVert\Sigma_{w}\rVert)^{2}.$ (103)
Compare the expression of $b_{s}(x_{0})$ and $\bar{C}$ in (85), we know that $b_{s}(x_{0})\leq\bar{C}$, and further $\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})\rVert\leq\bar{C}$. From (101), if
$\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\leq\min\{b_{s}(x_{0}),\bar{C}-b_{s}(x_{0})\}=:\epsilon_{v},$ (104)
then $\lVert\hat{C}_{\bar{K}_{k}}-\hat{b}_{s}(x_{0})\rVert\leq\bar{C}$. Therefore, the required sample for variance reduction is at least not larger than the case without variance reduction. Following the proof of Theorem 4, we can obtain the expression of $N_{3}$:
$N_{3}:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{11}(C(K),\delta_{x},\hat{b}_{s}(x_{0}))\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right],$ (105)
where
$\alpha_{11}(C(K),\delta_{x},\hat{b}_{s}(x_{0})):=\max\{n_{x},n_{u}\}^{2}\alpha^{2}_{12}(\hat{b}_{s}(x_{0}))+\left(\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))\right)^{2}.$
and $\alpha_{12}(\hat{b}_{s}(x_{0})):=\frac{n_{x}n_{u}}{r}\{\max\{\bar{C}-{b}_{s}(x_{0}),{b}_{s}(x_{0})\}+\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\}$.
From Algorithm 2 and Lemma 9, similar to the proof of Theorem 4, we have:
$\hat{C}_{k}^{V}\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/l}}\right)^{2}=:\bar{C}^{V}(\tilde{\delta}_{x})$ (106)
with predefined probability $\delta_{x}^{V}$. The expectation of $\hat{C}_{k}^{V}$ is upper bounded as:
$\mathbb{E}[\hat{C}_{k}^{V}]\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\lVert\Sigma_{w}\rVert\right)^{2}=:\bar{C}_{E}^{V}.$ (107)
The bound on $\lVert C^{V}_{k}-\mathbb{E}[\hat{C}_{k}^{V}]\rVert$ is given by:
$\lVert C^{V}_{k}-\mathbb{E}[\hat{C}_{k}^{V}]\rVert\leq\bar{C}_{E}^{V}+\bar{C}^{V}(\tilde{\delta}_{x})$ (108)
Similarly, the bound on $\lVert\mathbb{E}(C^{V}_{k}C^{V~\top}_{k})-{C}_{k}^{V}{C}_{k}^{V~\top}\rVert$ is given by:
$\begin{split}&\lVert\mathbb{E}(C^{V}_{k}C^{V~\top}_{k})-{C}_{k}^{V}{C}_{k}^{V~\top}\rVert\leq\max_{C^{V}_{k}}(\lVert C^{V}_{k}\rVert_{F})^{2}+\lVert\mathbb{E}[C^{V}_{k}]\rVert^{2}_{F}\leq(\bar{C}_{E}^{V})^{2}+(\bar{C}^{V}(\tilde{\delta}_{x}))^{2}.\end{split}$ (109)
Using the matrix concentration inequality [10, Lemma B2.5], we get:
$\tilde{n}_{\min}=\frac{2}{(\epsilon^{\prime}_{v})^{2}}\left((\bar{C}_{E}^{V}+\bar{C}^{V})^{2}+\frac{((\bar{C}_{E}^{V})^{2}+(\bar{C}^{V}(\tilde{\delta}_{x}))^{2})\epsilon_{v}}{3}\right)\log\left[\frac{2}{(\tilde{\delta}_{v})}\right].$ (110)
Thus, if the number of rollouts used to estimate the baseline function satisfies $n_{v}\geq\tilde{n}_{\min}(\tilde{\delta}_{x},\delta_{v})$, the sample size required to estimate the gradient with the baseline function is at least not larger than the case without the baseline function, with probability $\delta_{v}=1-(1-\tilde{\delta}_{v})(1-\tilde{\delta}_{x})$.
C.8 Proof of Theorem 7
Proof 14
Let $K_{i+1}:=\hat{K}_{i}-\eta_{i}\nabla C(\hat{K}_{i})$, where $\nabla C(\hat{K}_{i})$ is the exact gradient. Our goal is to have
$|C(\hat{K}_{i+1})-C(K_{i+1})|\leq\frac{\eta_{i}\sigma\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{2\lVert\Sigma_{K}^{*}\rVert}\epsilon$
Then when $C(\hat{K}_{i})-C(K^{*})\geq\epsilon$ and with (15), we have
$C(\hat{K}_{i+1})-C(K^{*})\leq\left(1-(1-\sigma)\frac{\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{2\lVert\Sigma_{K}^{*}\rVert}\right)(C(\hat{K}_{i})-C(K^{*}))$
Then convergence is guaranteed. Observing that: $\hat{K}_{i+1}-K_{i+1}=\eta_{i}(\hat{\nabla}C(\hat{K}_{i})-\nabla C(K_{i}))$, together with Theorem 1:
$\begin{split}|C(\hat{K}_{i+1})-C(K_{i+1})|&\leq h_{\nabla}(C(\hat{K}_{i}))\lVert\hat{K}_{i+1}-K_{i+1}\rVert\\ &=\eta_{i}h_{\nabla}(C(\hat{K}_{i}))\lVert\hat{\nabla}C(\hat{K}_{i})-\nabla C(K_{i})\rVert.\end{split}$
Then we can derive the requirement on the estimation accuracy $\epsilon_{\mathrm{PGD}}$ on $\hat{\nabla}C(\hat{K}_{i})$
$\epsilon_{\mathrm{PGD}}:=\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{2h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}.$
Then we should choose the $l,n,r$ as discussed in Theorem 4 to satisfy the accuracy and probability $\delta$ above. The gradient descent algorithm can reach the desired accuracy $\epsilon$ with probability $1-\delta$. The remainder of the proof follows Theorem 1, which assumes that the gradient estimate meets the required accuracy at each iteration, guaranteed by the union bound and using the fact $(1-\delta)^{n_{\mathrm{PGD}}}\geq 1-\delta n_{\mathrm{PGD}}$.
C.9 Proof of Theorem 8
Proof 15
Let ${K}_{i+1}:=\hat{K}_{i}-\eta_{i}{\nabla}C(\hat{K}_{i}){\Sigma}_{\hat{K}_{i}}^{-1}$. From the model-based natural policy gradient in Theorem 2:
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{\eta_{i}\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{K}^{*}\rVert}\right)(C(\hat{K}_{i})-C(K^{*}))$
we aim to bound $\lVert C(\hat{K}_{i+1})-C({K}_{i+1})\rVert\leq\frac{\epsilon\sigma\eta_{i}\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{2\lVert\Sigma_{K}^{*}\rVert}$, then when $C(K)-C(K^{*})\geq\epsilon$, we have
$C(\hat{K}_{i+1})-C(K^{*})\leq\left(1-(1-\sigma)\frac{\eta_{i}\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{2\lVert\Sigma_{K}^{*}\rVert}\right)(C(\hat{K}_{i})-C(K^{*}))$
We aim to have the following, for any $\hat{K}_{i}$:
$\lVert\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\nabla}C(\hat{K}_{i}){\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{2h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}.$
This is broken into two terms:
$\lVert\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\nabla}C(\hat{K}_{i}){\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert+\lVert{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\Sigma}_{\hat{K}_{i}}^{-1}\rVert$
From Weyl’s Theorem [9, Theorem 31], if $\epsilon\leq\frac{\lambda_{1}(\Sigma_{w})}{2}$, then it holds that $\lambda_{1}(\hat{\Sigma}_{\hat{K}_{i}})\leq\lambda_{1}({\Sigma}_{w})-\lVert\hat{\Sigma}_{\hat{K}_{i}}-\Sigma_{\hat{K}_{i}}\rVert=\frac{\lambda_{1}({\Sigma}_{w})}{2}$. Because $\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{2}{\lambda_{1}(\Sigma_{w})}$, we aim to make sure that
$\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}=:\epsilon_{NPG}^{C},$
so that $\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}$. Then we should choose the $l_{1},n_{1},r_{1}$ in Algorithm 1 according to Theorem 4 to satisfy the accuracy on $\hat{\nabla}C(\hat{K}_{i})$ and the probability $\delta^{\frac{1}{2}}$ above.
For the second term, $\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\Sigma}_{\hat{K}_{i}}^{-1}\rVert$ must be bounded by:
$\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K^{*}}\rVert\lVert{\nabla}C(\hat{K}_{i})\rVert}.$
where $h_{\nabla}$ is defined in (59). By matrix perturbation, if $\lambda_{1}(\Sigma_{\hat{K}_{i}})\geq\lambda_{1}(\Sigma_{w})$ and $\lVert\hat{\Sigma}_{\hat{K}_{i}}-{\Sigma}_{\hat{K}_{i}}\rVert\leq\frac{\lambda_{1}(\Sigma_{w})}{2}$, then $\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{2\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\Sigma}_{\hat{K}_{i}}^{-1}\rVert}{\lambda_{1}^{2}(\Sigma_{w})}$. The desired accuracy on $\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\Sigma}_{\hat{K}_{i}}^{-1}\rVert$ can be derived as:
$\begin{split}\lVert\hat{\Sigma}_{\hat{K}_{i}}-{\Sigma}_{\hat{K}_{i}}\rVert&\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\lVert{\nabla}C(\hat{K}_{i})\rVert}\\ &\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\sqrt{b_{\nabla}(C(\hat{K}_{i}))}}=:\epsilon_{NPG}^{\Sigma}\end{split}$ (111)
Then we should choose the $l_{2},n_{2},r_{2}$ according to Theorem 5 to satisfy the accuracy on $\hat{\Sigma}_{K}$ and probability $\delta^{\frac{1}{2}}$ above. If we choose $n\geq\max\{n_{1},n_{2}\},l\geq\max\{l_{1},l_{2}\},R\leq\min\{r_{1},r_{2}\}$, then we have
$\mathbb{P}\left\{\lVert C(\hat{K}_{i+1})-C({K}_{i+1})\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{2\lVert\Sigma_{K}^{*}\rVert}\right\}\geq 1-\delta$
Then natural policy gradient can reach the desired accuracy $\epsilon$ with probability $1-\delta$. The remainder of the proof follows Theorem 2, which assumes that the gradient estimate and covariance matrix meet the required accuracy at each iteration, guaranteed by the union bound and using the fact $(1-\delta)^{n_{\mathrm{NPG}}}\geq 1-\delta n_{\mathrm{NPG}}$.